### v5

# Decentralized Parametric Stability Certificates for Grid-Forming Converter Control

Abstract
We propose a decentralized framework for guaranteeing the small-signal stability of future power systems with grid-forming converters. Our approach leverages dynamic loop-shifting techniques to compensate for the lack of passivity in the network dynamics and establishes decentralized parametric stability certificates, depending on the local device-level controls and incorporating the effects of the network dynamics. By following practical tuning rules, we are able to ensure plug-and-play operation without centralized coordination. Unlike prior works, our approach accommodates coupled frequency and voltage dynamics, incorporates network dynamics, and does not rely on specific network configurations or operating points, offering a general and scalable solution for the integration of power-electronics-based devices into future power systems. We validate our theoretical stability results through numerical case studies in a high-fidelity simulation model.
{IEEEkeywords}
power system stability, grid-forming converter, passivity, decentralized stability conditions
1 Introduction
[…]
The transition to future power systems is characterized by a substantial increase in the share of power electronics (PE)-based generation devices. This shift introduces significant changes in system dynamics, where the interactions between PE-based generation and the remainder of the power network are not fully understood yet [1]. In particular, the fast and complex dynamics of PE-based devices can lead to unexpected effects (e.g., overvoltages [2], subsynchronous oscillations [3], etc.), posing considerable challenges for maintaining system stability. Consequently, there is a pressing need for a deeper understanding of how PE-based devices interact with one another or with the grid and for the development of robust stability frameworks to ensure stable operation.
One of the key aspects of future power systems involves the integration of PE-based devices with either grid-following or grid-forming control. Grid-forming (GFM) control is particularly promising as it enhances stability by establishing a well-defined ac voltage at the grid connection [1, 4]. However, it has been observed that network dynamics can significantly influence the stability of PE-dominated systems [5, 6], unlike traditional synchronous generator-dominated grids, which exhibit relatively slow dynamics compared to the fast network dynamics. This creates a need for analysis methods that account for the interplay between network dynamics and the control characteristics of PE-based devices.
In this paper, we address the destabilizing effects of network dynamics on the small-signal stability of heterogeneous interconnected GFM converters. Our approach seeks to compensate for these effects by proposing parametric decentralized stability certificates to ensure system-wide stability. Specifically, we introduce a decentralized framework for quantitatively certifying stability at the individual device level through local tuning rules for each GFM controller. This enables plug-and-play operation without centralized coordination. Our approach leverages dynamic loop-shifting techniques to compensate for the lack of passivity in the network dynamics. By formulating device-level specifications that are sufficiently passive, we ensure overall system stability in a flexible and scalable way.
Our contribution improves significantly over prior works on stability certification in GFM converter systems. Unlike [7, 8], which are limited to single-input single-output (SISO) dynamics, our approach accommodates multiple-input multiple-output (MIMO) dynamics coupling frequency and voltage, thus providing a more comprehensive analysis of the system behavior. Additionally, we incorporate the network dynamics, extending beyond the quasi-stationary or zero-power flow approximations in [7, 9, 10]. Our decentralized stability conditions enable the use of heterogeneous GFM models, improving over the homogeneity assumptions in [8, 11]. Moreover, unlike [12], our approach does not require detailed knowledge of the network configuration. Finally, our stability conditions are independent of the operating point, provided small bus angle differences, unlike [9], where the the stability conditions depend on specific operating points. Collectively, these attributes allow our framework to provide, to the best of our knowledge, the most general, explicitly parametric, and practically applicable small-signal stability certificates for interconnected GFM devices in the literature. Since our stability conditions are readily actionable for control tuning and grid-code design, they can ensure a stable and scalable integration of PE-based generation in future power systems.
The paper is structured as follows. Section 2 introduces preliminary concepts of feedback stability and passivity for linear time-invariant (LTI) systems. In Section 3, we describe the dynamic power system model utilized for stability analysis. Section 4 presents our main results: decentralized parametric stability certificates for GFM converters that account for network dynamics. The results are validated through numerical case studies in Section 5. Section 6 concludes the paper.
2 Foundations of Interconnected Systems
2.1 Preliminaries
Let $\mathbb{R}$ denote the set of real numbers, $\mathbb{N}$ the set of positive integers, and $\mathbb{C}$ the set of complex numbers with imaginary unit $\mathrm{j}$. We use $I_{n}$ to denote the $n$-by-$n$ identity matrix (abbreviated as $I$ when the dimensions can be inferred from the context). We use $\otimes$ to denote the Kronecker product. We use $A=\mathrm{diag}(A_{1},A_{2},...,A_{k}),\,k\in\mathbb{Z}$ to denote the block-diagonal matrix with blocks $A_{1},A_{2},...,A_{k}$. The Euclidean norm of a matrix $A$ is defined as $||A||_{F}=\sqrt{\textstyle\sum_{i,j}|a_{ij}|^{2}}$, where $a_{ij}$ is the entry in the $i$th row and $j$th column of $A$.
For a complex matrix $A\in\mathbb{C}^{n\times n}$, we use $A^{\star}$ to denote its conjugate transpose. A matrix $A\in\mathbb{C}^{n\times n}$ satisfying $A=A^{\star}$ is called Hermitian. A Hermitian matrix $A\in\mathbb{C}^{n\times n}$ is said to be positive definite (semi-definite), denoted by $A\succ 0$ $(\succeq 0)$, if $x^{\star}Ax>0$ $(\geq 0)$ for all $x\neq 0$. A matrix $A\in\mathbb{C}^{n\times n}$ is called diagonally dominant, if $|a_{ii}|\geq\textstyle\sum_{j\neq i}|a_{ij}|,\,\forall i$.
To establish our main stability results in Section 4, we require the following lemma which follows from the Gershgorin’s Circle Theorem [13, Thm. 6.1.10]:
Lemma 1.
A Hermitian diagonally dominant matrix with real non-negative diagonal entries is positive semi-definite.
Further, we review the recent concept of phases of complex matrices based on the matrix’s numerical range [14]. Namely, the numerical range of a complex matrix $A\in\mathbb{C}^{n\times n}$ is defined as $W(A)=\{x^{\star}Ax:x\in\mathbb{C}^{n},||x||=1\}$. If $0\notin W(A)$, then $A$ is said to be a sectorial matrix. For a sectorial $A$, there exists a nonsingular matrix $T$ and a diagonal unitary matrix $D$ such that $A=T^{\star}DT$ [14]. The diagonal elements of $D$ are distributed in an arc on the unit circle with length smaller than $\pi$. Then, the phases of $A$, denoted by
$\overline{\phi}(A)=\phi_{1}(A)\geq\dots\geq\phi_{n}(A)=\underline{\phi}(A),$ (1)
are defined as the phases of the diagonal entries of $D$ so that $\overline{\phi}(A)-\underline{\phi}(A)<\pi$. The definition of phases can be generalized to so-called semi-sectorial matrices whose numerical ranges contain the origin on their boundaries and $\overline{\phi}(A)-\underline{\phi}(A)\leq\pi$.
2.2 Transfer Functions & Stability
Let $u(t)\in\mathbb{R}^{n}$ be the input and $y(t)\in\mathbb{R}^{n}$ the output signal of a MIMO LTI system. The transfer matrix $H(s)\in\mathbb{C}^{n\times n}$ describes the input-output system response in the frequency domain as $y(s)=H(s)u(s)$, where $y(s)$ and $u(s)$ are the Laplace transformations of the output and input, respectively.
We denote the space of $n$-by-$n$ real rational proper transfer matrices of stable LTI systems by $\mathcal{RH}_{\infty}^{n\times n}$. An LTI system is called stable if all poles are in the open left half plane (LHP). It is called semi-stable if it may have poles on the imaginary axis but no poles in the open right half plane (RHP).
Definition 1 (Internal Feedback Stability [15]).
Let $H_{1}(s)$ and $H_{2}(s)$ be $n\times n$ real rational proper transfer matrices. The feedback system in Fig. 1(a) is internally feedback stable if and only if the following four closed-loop transfer matrices
$\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}=\underset{\eqqcolon\,H_{1}\#H_{2}(s)}{\underbrace{\begin{bmatrix}(I+H_{1}H_{2})^{-1}H_{1}&-(I+H_{1}H_{2})^{-1}H_{1}H_{2}\\ H_{2}(I+H_{1}H_{2})^{-1}H_{1}&H_{2}(I+H_{1}H_{2})^{-1}\end{bmatrix}}}\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}$ (2)
are stable, compactly referred to as $H_{1}\#H_{2}\in\mathcal{RH}_{\infty}^{2n\times 2n}$.
If there are no RHP pole-zero cancellations between $H_{1}(s)$ and $H_{2}(s)$, then stability of one closed-loop transfer matrix implies stability of the others [15, Thm 4.7]:
Lemma 2.
Assume there are no RHP pole-zero cancellations between $H_{1}(s)$ and $H_{2}(s)$, i.e., all RHP poles in $H_{1}(s)$ and $H_{2}(s)$ are contained in the minimal realization of $H_{1}(s)H_{2}(s)$ and $H_{2}(s)H_{1}(s)$. Then, the feedback system in Fig. 1(a) is internally feedback stable if and only if one of the four closed-loop transfer matrices in 2 is stable.
$H_{1}(s)$$H_{2}(s)$--$w_{1}$$y_{1}$$y_{2}$$w_{2}$
(a) Original feedback system.
[…]
(b) Loop shifting with $\Gamma(s)$.
Figure 1: Closed-loop feedback interconnection of two LTI systems.
2.3 Small-Phase & Passivity Theory
For LTI systems, the property of passivity is equivalent to the property of positive realness [14, 16, 17, 18].
Definition 2 (Passive Transfer Matrix [14]).
A $n\times n$ real rational proper transfer matrix $H(s)$ is passive if
  1. (i)
    poles of all elements of $H(s)$ are in $\text{Re}(s)\leq 0$,
  2. (ii)
    $H(\mathrm{j}\omega)+H^{\star}(\mathrm{j}\omega)\succeq 0$ for any $\omega$ for which $\mathrm{j}\omega$ is not a pole of any element of $H(s)$,
  3. (iii)
    any purely imaginary pole $\mathrm{j}\omega$ of any element of $H(s)$ is a simple pole and $\text{lim}_{s\rightarrow\mathrm{j}\omega}(s-\mathrm{j}\omega)H(s)$ is positive semi-definite Hermitian.
Definition 3 (Strictly Passive Transfer Matrix [14]).
A $n\times n$ real rational proper transfer matrix $H(s)$ is strictly passive if
  1. (i)
    poles of all elements of $H(s)$ are in $\text{Re}(s)<0$,
  2. (ii)
    $H(\mathrm{j}\omega)+H^{\star}(\mathrm{j}\omega)\succ 0$ for any $\omega\in[-\infty,\infty]$.
We now present a generalization of the passivity theorem [17, 18] for the special case of LTI systems as follows:
Theorem 1 (Passivity Theorem).
Consider two LTI systems $H_{1}(s)$ and $H_{2}(s)$ in negative feedback configuration, as shown in Fig. 1(a). The feedback system is internally feedback stable if $H_{1}(s)$ is strictly passive and $H_{2}(s)$ is passive.
The above theorem is widely known in the SISO case [17], but we could not locate a proof in this general MIMO setting. We thus provide a self-contained proof below, relying on the Generalized Small-Phase Theorem [14, Thm 7.1]:
Lemma 3 (Generalized Small-Phase Theorem).
Let $H_{1}(s)\in\mathcal{RH}_{\infty}^{n\times n}$ be frequency-wise sectorial and $H_{2}(s)$ be semi-stable frequency-wise semi-sectorial with $\mathrm{j}\Omega$ being the set of poles on the imaginary axis. Then, $H_{1}\#H_{2}$ in Fig. 1(a) is internally feedback stable if
$\begin{split}&\underset{\omega\in[0,\infty]\backslash\Omega}{\mathrm{sup}}\left[\overline{\phi}(H_{1}(\mathrm{j}\omega))+\overline{\phi}(H_{2}(\mathrm{j}\omega))\right]<\pi\quad\text{and}\\ &\underset{\omega\in[0,\infty]\backslash\Omega}{\mathrm{inf}}\left[\underline{\phi}(H_{1}(\mathrm{j}\omega))+\underline{\phi}(H_{2}(\mathrm{j}\omega))\right]>-\pi.\end{split}$ (3)
Proof of Theorem 1.
We prove Theorem 1 by applying Lemma 3 as follows: Since $H_{1}(s)$ is strictly passive, it follows that $H_{1}(s)\in\mathcal{RH}_{\infty}^{n\times n}$ is frequency-wise sectorial with phases $\left[\underline{\phi}(H_{1}(\mathrm{j}\omega)),\overline{\phi}(H_{1}(\mathrm{j}\omega))\right]\subset(-\tfrac{\pi}{2},\tfrac{\pi}{2}),\forall\omega\in[0,\infty]$ (see [14] for a detailed derivation). Likewise, since $H_{2}(s)$ is passive, it follows that $H_{2}(s)$ is semi-stable with poles at $\mathrm{j}\Omega$ on the imaginary axis and frequency-wise semi-sectorial with phases $\left[\underline{\phi}(H_{2}(\mathrm{j}\omega)),\overline{\phi}(H_{2}(\mathrm{j}\omega))\right]\subset[-\tfrac{\pi}{2},\tfrac{\pi}{2}],\forall\omega\in[0,\infty]\backslash\Omega$ (see [14] for a detailed derivation). Therefore, the conditions in 3 are satisfied which implies $H_{1}\#H_{2}\in\mathcal{RH}_{\infty}^{2n\times 2n}$. ∎
To extend the passivity-based stability conditions to more general feedback interconnections with both passive and nonpassive subsystems, we employ the concepts of excess and shortage of passivity [17, 18]. The basic idea is that the excess of passivity of one subsystem can compensate for the passivity deficit in the other subsystem, such that their feedback interconnection remains stable. This can be achieved by performing loop-shifting techniques as shown in Fig. 1, where a transfer matrix $\Gamma(s)\in\mathbb{C}^{n\times n}$ is added as a positive feedforward to $H_{2}(s)$, and as a positive feedback to $H_{1}(s)$. Now the idea is that if the original subsystems $H_{1}(s)$ and $H_{2}(s)$ in Fig. 1(a) are not satisfying the passivity conditions in Theorem 1, the two subsystems ${H}^{\prime}_{2}(s)$ and ${H}^{\prime}_{1}(s)$ in Fig. 1(b) might do so for a suitable $\Gamma(s)$, thus resulting in a stable feedback interconnection ${H}^{\prime}_{1}\#{H}^{\prime}_{2}$ (for details see literature on frequency-dependent passivity indices, e.g., [17, 18]).
Among the four closed-loop transfer matrices in 2 relevant to internal feedback stability, it is immediately evident that only the upper-left transfer matrix mapping from $w_{1}$ to $y_{1}$ remains equivalent between the original system in Fig. 1(a) and the loop-shifted system in Fig. 1(b). Specifically, $(I+{H}^{\prime}_{1}{H}^{\prime}_{2})^{-1}{H}^{\prime}_{1}=(I+H_{1}H_{2})^{-1}H_{1}$. Consequently, stability of ${H}^{\prime}_{1}\#{H}^{\prime}_{2}$ directly implies stability of $(I+H_{1}H_{2})^{-1}H_{1}$. If there are no RHP pole-zero cancellations between $H_{1}(s)$ and $H_{2}(s)$, by Lemma 2, stability of the upper-left transfer matrix $(I+H_{1}H_{2})^{-1}H_{1}$ also guarantees stability of $H_{1}\#H_{2}$.
Remark 1.
By swapping the feedforward and feedback of $\Gamma(s)$ in Fig. 1(b) and comparing the closed-loop transfer function from $w_{2}$ to $y_{2}$, similar stability conclusions can be made.
3 Power System Model
3.1 Small-Signal Network Dynamics
We study the stabilization of multiple three-phase GFM voltage source converters (VSCs) connected through a dynamic transmission network modeled by resistive-inductive lines (Fig. 2). We consider a Kron-reduced [19], balanced network with $n\in\mathbb{N}$ converter nodes, denoted by $\{1,...,n\}$, where the dynamic small-signal model (in the global per unit system) of the line between node $i$ and node $j$ ($i,j\in\{1,...,n\}$) is given in the frequency domain as
$\begin{bmatrix}\Delta i_{\mathrm{d},ij}(s)\\ \Delta i_{\mathrm{q},ij}(s)\end{bmatrix}=y_{ij}(s)\left(\begin{bmatrix}\Delta v_{\mathrm{d},i}(s)\\ \Delta v_{\mathrm{q},i}(s)\end{bmatrix}-\begin{bmatrix}\Delta v_{\mathrm{d},j}(s)\\ \Delta v_{\mathrm{q},j}(s)\end{bmatrix}\right),$ (4)
where $\Delta i_{\mathrm{dq},ij}=\left[\Delta i_{\mathrm{d},ij}\,\,\Delta i_{\mathrm{q},ij}\right]^{\top}$ is the current vector from node $i$ to node $j$ in the global $\mathrm{dq}$ coordinates with a constant nominal rotating frequency $\omega_{0}$, $\Delta v_{\mathrm{dq},i}=\left[\Delta v_{\mathrm{d},i}\,\,\Delta v_{\mathrm{q},i}\right]^{\top}$ is the voltage vector of node $i$, and $y_{ij}(s)$ is a $2\times 2$ transfer matrix encoding the small-signal dynamics of the line, i.e.,
$y_{ij}(s)=b_{ij}\underset{\eqqcolon\,f_{\rho}(s)}{\underbrace{\begin{bmatrix}\rho+\tfrac{s}{\omega_{0}}&1\\ -1&\rho+\tfrac{s}{\omega_{0}}\end{bmatrix}\tfrac{1}{1+\left(\rho+\tfrac{s}{\omega_{0}}\right)^{2}}}},$ (5)
where $b_{ij}=\tfrac{1}{l_{ij}}$ is the line susceptance, and $\rho_{ij}=\tfrac{r_{ij}}{l_{ij}}$ is the resistance-inductance ratio of the line $ij$ which is assumed to be small and uniform (i.e., $\rho_{ij}=\rho\ll 1,\,\forall\,i,j\in\{1,...,n\}$) throughout the dominantly inductive transmission network. Notice that $b_{ij}=0$ and $\rho_{ij}=0$ if there is no line.
[…]
Figure 2: Illustration of the multi-converter transmission system.
The bus current injection of each node $i$ is defined as $\Delta i_{\mathrm{dq},i}(s)\coloneqq\textstyle\sum_{j\neq i}^{n}\Delta i_{\mathrm{dq},ij}(s)$, based on which we can construct the network dynamics for all nodes in the form of the $2n\times 2n$ Laplacian transfer matrix $Y(s)$, i.e.,
$\begin{bmatrix}\Delta i_{\mathrm{d},1}(s)\\ \Delta i_{\mathrm{q},1}(s)\\ \vdots\\ \Delta i_{\mathrm{d},n}(s)\\ \Delta i_{\mathrm{q},n}(s)\end{bmatrix}=\underset{\eqqcolon\ Y(s)}{\underbrace{B\otimes f_{\rho}(s)}}\begin{bmatrix}\Delta v_{\mathrm{d},1}(s)\\ \Delta v_{\mathrm{q},1}(s)\\ \vdots\\ \Delta v_{\mathrm{d},n}(s)\\ \Delta v_{\mathrm{q},n}(s)\end{bmatrix},$ (6)
where $Y_{ij}(s)=-y_{ij}(s)$ if $i\neq j$, $Y_{ii}(s)=\textstyle\sum_{j\neq i}^{n}y_{ij}(s)$, and $B\in\mathbb{R}^{n\times n}$ is a Laplacian matrix
$B=\begin{bmatrix}\textstyle\sum_{j\neq 1}^{n}b_{1j}&\dots&-b_{1n}\\ \vdots&\ddots&\vdots\\ -b_{n1}&\cdots&\textstyle\sum_{j\neq n}^{n}b_{nj}\end{bmatrix},$ (7)
that encodes the network topology and line susceptances, where $B_{ij}=-b_{ij}$ if $i\neq j$, and $B_{ii}=\textstyle\sum_{j\neq i}^{n}b_{ij}$.
The time-domain bus voltages are modeled in polar coordinates as $v_{\mathrm{d},i}(t)\coloneqq|v|_{i}(t)\cos{\delta_{i}(t)}$ and $v_{\mathrm{q},i}(t)\coloneqq|v|_{i}(t)\sin{\delta_{i}(t)}$ with magnitude $|v|_{i}(t)$ and relative angle $\delta_{i}(t)=\theta_{i}(t)-\theta_{0}(t)$ in SI units, where $\theta_{i}(t)$ is the voltage angle at bus $i\in\{1,...,n\}$, and $\theta_{0}(t)$ the Park transformation angle of the global $\mathrm{dq}$ coordinates, which can be expressed as
$\begin{split}|v|_{i}(t)=\sqrt{v_{\mathrm{d},i}(t)^{2}+v_{\mathrm{q},i}(t)^{2}},\quad\delta_{i}(t)=\arctan\left(\tfrac{v_{\mathrm{q},i}(t)}{v_{\mathrm{d},i}(t)}\right).\end{split}$ (8)
We consider the active and reactive branch powers
$\begin{split}p_{ij}(t)&=v_{\mathrm{d},i}(t)i_{\mathrm{d},ij}(t)+v_{\mathrm{q},i}(t)i_{\mathrm{q},ij}(t)\\ q_{ij}(t)&=-v_{\mathrm{d},i}(t)i_{\mathrm{q},ij}(t)+v_{\mathrm{q},i}(t)i_{\mathrm{d},ij}(t),\end{split}$ (9)
and define the associated bus power injections $p_{i}(t)\coloneqq\textstyle\sum_{j\neq i}^{n}p_{ij}(t)$ and $q_{i}(t)\coloneqq\textstyle\sum_{j\neq i}^{n}q_{ij}(t)$, which are obtained as
$\begin{split}p_{i}(t)&=v_{\mathrm{d},i}(t)i_{\mathrm{d},i}(t)+v_{\mathrm{q},i}(t)i_{\mathrm{q},i}(t)\\ q_{i}(t)&=-v_{\mathrm{d},i}(t)i_{\mathrm{q},i}(t)+v_{\mathrm{q},i}(t)i_{\mathrm{d},i}(t).\end{split}$ (10)
By linearizing 8 and 10, transforming them into the frequency domain, and performing some analytical computations (see Appendix 7 for details), we obtain
$\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta p(s)\\ \Delta q(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta p_{1}(s)\\ \Delta q_{1}(s)\\ \vdots\\ \Delta p_{n}(s)\\ \Delta q_{n}(s)\end{bmatrix}}}=\underset{\eqqcolon\,{N}(s)}{\underbrace{\begin{bmatrix}N_{11}(s)&\cdots&N_{1n}(s)\\ \vdots&\ddots&\vdots\\ N_{n1}(s)&\cdots&N_{nn}(s)\end{bmatrix}}}\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta\delta(s)\\ \Delta|v|_{\mathrm{n}}(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta\delta_{1}(s)\\ \Delta{|v|}_{\mathrm{n},1}(s)\\ \vdots\\ \Delta\delta_{n}(s)\\ \Delta{|v|}_{\mathrm{n},n}(s)\end{bmatrix}}}$ (11)
with the $2\times 2$ transfer matrix blocks $N_{ii}(s)$ and $N_{ij}(s)$
$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}1&\rho+\tfrac{s}{\omega_{0}}\\ -(\rho+\tfrac{s}{\omega_{0}})&1\end{bmatrix}\\ &\quad\quad\quad+\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}-1&\rho\\ \rho&1\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}-1&-(\rho+\tfrac{s}{\omega_{0}})\\ (\rho+\tfrac{s}{\omega_{0}})&-1\end{bmatrix},\end{split}$ (12)
where $\Delta{|v|}_{\mathrm{n},i}(s)\coloneqq\tfrac{\Delta|v|_{i}(s)}{|v|_{0,i}}$ is the voltage magnitude at bus $i$ normalized at the steady-state $|v|_{0,i}$. Moreover, to derive 12, as in any power system small-signal model, we have assumed a small steady-state angle difference $\delta_{0,i}\approx\delta_{0,j}$, which is standard and justified since thermal limitations for transmission lines preclude large angle differences [20].
For a dominantly inductive transmission network with $\rho\ll 1$, the standard approximation of the line dynamics matrix
$\begin{split}M(s)&=\tfrac{1}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}1&\rho+\tfrac{s}{\omega_{0}}\\ -(\rho+\tfrac{s}{\omega_{0}})&1\end{bmatrix}\end{split}$ (13)
appearing in $N_{ii}(s)$ and $N_{ij}(s)$ in 12 is
$\begin{split}M_{1}(s)&=\tfrac{1}{1+\tfrac{s^{2}}{\omega_{0}^{2}}}\begin{bmatrix}1&\tfrac{s}{\omega_{0}}\\ -\tfrac{s}{\omega_{0}}&1\end{bmatrix},\end{split}$ (14)
i.e., losses are entirely neglected ($\rho=0$). Here we put forward a novel and more accurate approximation for $M(s)$, namely
$M_{2}(s)$ $=\tfrac{1}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}1&\tfrac{s}{\omega_{0}}\\ -\tfrac{s}{\omega_{0}}&1\end{bmatrix}.$ (15)
Indeed, a straightforward calculation comparing the Euclidean norms of the matrix distances for $s=\mathrm{j}\omega$ reveals that $\forall\omega\geq 0$
$\tfrac{||M(\mathrm{j}\omega)-M_{2}(\mathrm{j}\omega)||_{F}}{||M(\mathrm{j}\omega)-M_{1}(\mathrm{j}\omega)||_{F}}=\sqrt{\tfrac{1+\tfrac{\omega^{4}}{\omega_{0}^{4}}-2\tfrac{\omega^{2}}{\omega_{0}^{2}}}{1+\tfrac{\omega^{4}}{\omega_{0}^{4}}+6\tfrac{\omega^{2}}{\omega_{0}^{2}}+\rho^{2}\left(1+\tfrac{\omega^{2}}{\omega_{0}^{2}}\right)}}<1,$ (16)
that is, $M_{2}$ is a strictly better approximation for $\rho>0$. For $\rho\rightarrow 0$, it can be shown that both approximations are consistent in terms of recovering $M$ asymptotically. Using the approximation in 15, we can therefore eliminate the less dominant terms of the off-diagonal matrix elements in 12 as stated below.
Definition 4 (Dynamic Small-Signal Network Model).
The transfer matrix blocks for the dynamic (i.e., $s\neq 0$, $|v|_{0,i}\neq|v|_{0,j}$) small-signal network model in 11 used to derive the stability certificates in Section 4 are given by:
$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}1&\tfrac{s}{\omega_{0}}\\ -\tfrac{s}{\omega_{0}}&1\end{bmatrix}\\ &\quad\quad\quad+\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}-1&-\tfrac{s}{\omega_{0}}\\ \tfrac{s}{\omega_{0}}&-1\end{bmatrix}.\end{split}$ (17)
By setting $s=0$ in 17, we derive the quasi-stationary network model, which is employed in a similar form in [9]:
Definition 5 (Network-Simplification Level 1).
The transfer matrix blocks for the quasi-stationary (i.e., $s=0$, $|v|_{0,i}\neq|v|_{0,j}$) small-signal network model in 11 are given by:
$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{1}{1+\rho^{2}}\begin{bmatrix}|v|_{0,i}|v|_{0,j}&0\\ 0&2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}-1&0\\ 0&-1\end{bmatrix}.\end{split}$ (18)
To further simplify 18, we set $|v|_{0,i}=|v|_{0,j}=|v|_{0}$, yielding the zero-power flow network model similar to [10]:
Definition 6 (Network-Simplification Level 2).
The transfer matrix blocks for the zero-power flow (i.e., $s=0$, $|v|_{0,i}=|v|_{0,j}=|v|_{0}$) small-signal network model in 11 are given by:
$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0}^{2}}{1+\rho^{2}}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0}^{2}}{1+\rho^{2}}\begin{bmatrix}-1&0\\ 0&-1\end{bmatrix}.\end{split}$ (19)
[…]
Figure 3: Basic implementation of a three-phase GFM VSC.
Remark 2.
For $\rho=0$, the network model in [9] corresponds to 18 in Definition 5, while the model in [10] aligns with 19 in Definition 6. Moreover, other works [8, 7] focus solely on SISO frequency dynamics. Hence, our dynamic network model in 17 offers the most detailed representation in literature which is compatible with theoretical stability certificates.
3.2 Grid-Forming Converter Dynamics
Fig. 3 shows the basic implementation of a GFM three-phase VSC connected to the power network in Fig. 2. The linearized small-signal dynamics (in the global per unit system) of the $i$-th VSC ($i\in\{1,...,n\}$) can be represented by the $2\times 2$ transfer matrix $D_{i}(s)$, mapping from the active and reactive power injections $\Delta p_{i}(s)$ and $\Delta q_{i}(s)$ to the imposed frequency and voltage magnitude $\Delta\omega_{i}(s)=\Delta\delta_{i}(s)s$ and $\Delta|v|_{i}(s)$, i.e.,
$-\begin{bmatrix}\Delta\omega_{i}(s)\\ \Delta|v|_{i}(s)\end{bmatrix}=D_{i}(s)\begin{bmatrix}\Delta p_{i}(s)\\ \Delta q_{i}(s)\end{bmatrix}.$ (20)
The most prevalent GFM control strategy is filtered droop control [21]. Assuming time-scale separation, we can neglect the internal dynamics of the VSC (see Appendix 8 for a detailed derivation) and consider $D_{i}(s)$ as in the following definition. Of course, all our developments will be evaluated on the full VSC model in Section 5.
Definition 7 (Dynamic Small-Signal Converter Model).
The small-signal converter dynamics 20 used to derive the stability certificates in Section 4 are given by the droop controller
$D_{i}(s)=\begin{bmatrix}\tfrac{d_{\mathrm{p},i}}{\tau_{\mathrm{p},i}s+1}&0\\ 0&\tfrac{d_{\mathrm{q},i}}{\tau_{\mathrm{q},i}s+1}\end{bmatrix},$ (21)
where $d_{\mathrm{p},i}\in\mathbb{R}_{>0}$ and $d_{\mathrm{q},i}\in\mathbb{R}_{>0}$ are the active and reactive power droop gains, and $\tau_{\mathrm{p},i}\in\mathbb{R}_{\geq 0}$ and $\tau_{\mathrm{q},i}\in\mathbb{R}_{\geq 0}$ are the active and reactive power low-pass filter time constants.
Finally, if we extend 20 to include the dynamics of all $n$ converters, we obtain the following $2n\times 2n$ transfer matrix
$-\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta\omega(s)\\ \Delta|v|(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta\omega_{1}(s)\\ \Delta{|v|}_{1}(s)\\ \vdots\\ \Delta\omega_{n}(s)\\ \Delta{|v|}_{n}(s)\end{bmatrix}}}=\underset{\eqqcolon\,D(s)}{\underbrace{\begin{bmatrix}D_{1}(s)&0_{2\times 2}&\dots&0_{2\times 2}\\ 0_{2\times 2}&D_{2}(s)&\dots&0_{2\times 2}\\ \vdots&\vdots&\ddots&\vdots\\ 0_{2\times 2}&0_{2\times 2}&\dots&D_{n}(s)\end{bmatrix}}}\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta p(s)\\ \Delta q(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta p_{1}(s)\\ \Delta q_{1}(s)\\ \vdots\\ \Delta p_{n}(s)\\ \Delta q_{n}(s)\end{bmatrix}}}.$ (22)
$\mathcal{D}_{0}(s)$$\mathcal{N}_{0}(s)$--$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|v|\end{bmatrix}$$\begin{bmatrix}\Delta p\\ \Delta q\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|v|_{\mathrm{d}}\end{bmatrix}$device dynamicsnetwork dynamics--$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$
Figure 4: Closed-loop power system dynamics where $\Delta p_{\mathrm{d}}$ and $\Delta q_{\mathrm{d}}$ are the active and reactive power disturbances, $\Delta\omega_{\mathrm{d}}$ is the frequency disturbance, and $\Delta|{v}|_{\mathrm{d}}$ is the voltage magnitude disturbance.
3.3 Dynamic Power System Model
The closed-loop power system dynamics are modeled as the feedback interconnection of the converter device dynamics in 22 and the network dynamics in 11 as illustrated in Fig. 4. We focus on the stability of the bus frequency and voltage magnitudes, given by $\Delta\omega_{i}(s)=\Delta\delta_{i}(s)s$ and $\Delta|v|_{i}(s)$ for $i\in\{1,...,n\}$ and consider these quantities as interconnection signals between the device and network dynamics in Fig. 4. Accordingly, the interconnected subsystems are defined as:
$\begin{split}\mathcal{D}_{0}(s)&\coloneqq D(s)\\ \mathcal{N}_{0}(s)&\coloneqq N(s)\cdot\mathrm{diag}(\tfrac{1}{s},\tfrac{1}{|v|_{0,1}},\dots,\tfrac{1}{s},\tfrac{1}{|v|_{0,n}}).\end{split}$ (23)
In the next section, we establish internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$ under certain decentralized parametric conditions.
$c_{i,\rho}$ Definition $c_{i,\rho}$ Definition
$c_{1,\rho}$ $({1+\rho^{2}})/({5+2\rho^{2}})$ $c_{8,\rho}$ $2\rho\omega_{0}(15|v|_{\mathrm{max}}^{2}+4)(10\rho^{2}|v|_{\mathrm{max}}^{2}+25|v|_{\mathrm{max}}^{2}+4)$
$c_{2,\rho}$ ${(1+\rho^{2})^{2}}/({6\rho})$ $c_{9,\rho}$ $10\omega_{0}\rho^{4}|v|_{\mathrm{max}}^{2}+95\omega_{0}\rho^{2}|v|_{\mathrm{max}}^{2}+16\omega_{0}\rho^{2}+25\omega_{0}|v|_{\mathrm{max}}^{2}$
$c_{3,\rho}$ $2\rho(1+\rho^{2})$ $c_{10,\rho}$ $5\omega_{0}(1+\rho^{2})(20\rho^{3}|v|_{\mathrm{max}}^{2}+80\rho|v|_{\mathrm{max}}^{2}+16\rho)$
$c_{4,\rho}$ $((2\rho^{2}-5)+\sqrt{(5-2\rho^{2})^{2}+16\rho^{2}})/4\rho$ $c_{11,\rho}$ $50\rho\omega_{0}(1+\rho^{2})^{2}$
$c_{5,\rho}$ ${(1+\rho^{2})}/({10\rho^{2}|v|_{\mathrm{max}}^{2}+25|v|_{\mathrm{max}}^{2}+4)}$ $c_{12,\rho}$ ${(\rho^{2}+1)^{2}}/({6\rho|v|_{\mathrm{max}}^{2}})$
$c_{6,\rho}$ ${5(1+\rho^{2})^{2}}/({30\rho|v|_{\mathrm{max}}^{2}+8\rho)}$ $c_{13,\rho}$ $-2\rho|v|_{\mathrm{max}}^{4}$
$c_{7,\rho}$ ${5(1+\rho^{2})}/{4}$ $c_{14,\rho}$ $|v|_{\mathrm{max}}^{2}(2\rho^{2}-5)$
Table 1: Definition of the quantities $c_{l,\rho}$ for $l=1,...,14$ in the conditions 24 and 25 depending on the parameter $\rho$.
(a) 2D plot with $\rho=0.3$.
(b) 3D plot with variable $\rho$.
Figure 5: Graphical illustration of the stability conditions in 24, where $\alpha_{\mathrm{p},i}=\tfrac{d_{\mathrm{p},i}}{\omega_{0}}|v|_{\mathrm{max}}^{2}\sum_{j\neq i}^{n}b_{ij}$ and $\tilde{\tau}_{\mathrm{p},i}=\tau_{\mathrm{p},i}\omega_{0}$,.
(a) 2D plot with $\rho=0.3$.
(b) 3D plot with variable $\rho$.
Figure 6: Graphical illustration of the stability conditions in 25, where $\alpha_{\mathrm{q},i}=\tfrac{d_{\mathrm{q},i}}{|v|_{0,i}}\sum_{j\neq i}^{n}b_{ij}$ and $\tilde{\tau}_{\mathrm{q},i}=\tau_{\mathrm{q},i}\omega_{0}$.
4 Decentralized Stability Conditions
By applying dynamic loop-shifting techniques and passivity theory (see Section 4.2 for details), we derive parametric decentralized stability conditions that serve as local tuning rules for the dynamic droop control 21 of each VSC $i\in\{1,\dots,n\}$, ensuring the internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$ in Fig. 4. The stability conditions for each VSC $i$ depend on
  • its own tunable local droop control parameters, namely, the droop gains $d_{\mathrm{p},i},\,d_{\mathrm{q},i}$ and the time constants $\tau_{\mathrm{p},i},\,\tau_{\mathrm{q},i}$,
  • certain network parameters, including the susceptances of the neighboring lines $b_{ij}$, the resistance-to-inductance ratio $\rho$, and the maximum steady-state bus voltage magnitude $|v|_{\mathrm{max}}$. These parameters are either locally accessible or globally agreed upon, e.g., in grid codes.
Crucially, the conditions are local and entirely independent of the control parameters of other VSCs, making them directly applicable for scalable and decentralized stability assessment, as well as for local controller tuning and grid-code design.
The decentralized stability conditions for each VSC $i$ can be categorized into decoupled constraints on the active power-frequency droop control and the reactive power-voltage droop control. For each control scheme, these conditions can be visualized in either a two-dimensional (for fixed $\rho$) or a three-dimensional (for variable $\rho$) coordinate system, as shown in Figures 5 and 6. The coordinate axes represent scaled versions of the local droop gains and time constants (in 2D) and the global network parameter $\rho$ (in 3D), with typical value ranges. The resulting feasible parameter sets are indicated by green dots. As illustrated in Fig. 5, for small values of $\rho$, closed-loop stability is ensured when the droop gain $d_{\mathrm{p},i}$ and/or the self-susceptance $\sum_{j\neq i}^{n}b_{ij}$ are sufficiently small, irrespective of the time constant $\tau_{\mathrm{p},i}$. Notably, this suggests potential instability issues in future transmission systems with a high density of devices interconnected by short lines, which correspond to large values of $\sum_{j\neq i}^{n}b_{ij}$. Further, interpreting the filter time constant $\tau_{\mathrm{p},i}$ as virtual inertia reveals that increasing virtual inertia (i.e., larger $\tau_{\mathrm{p},i}$) does not necessarily enhance system stability. These findings align with the small-signal stability conditions derived for active power droop control in conjunction with SISO frequency dynamics in [8].
Similarly, Fig. 6 shows that closed-loop stability is guaranteed when the droop gain $d_{\mathrm{q},i}$ and/or the self-susceptance $\sum_{j\neq i}^{n}b_{ij}$ are sufficiently small, provided that $\tau_{\mathrm{q},i}$ is nonzero. In particular, in contrast to $\tau_{\mathrm{p},i}$, a larger $\tau_{\mathrm{q},i}$ can be stabilizing.
Finally, from the 3D plots in Figures 5 and 6, it becomes apparent how an increasing $\rho$ allows for larger local droop gains of both the active and reactive power droop control.
In what follows, we provide an algebraic parametrization of the conditions in Figures 5 and 6 (a detailed derivation is provided in Section 4.2). Namely, for the active power-frequency droop control, we require
$\begin{split}\alpha_{\mathrm{p},i}&<c_{1\rho}(2\rho+\tilde{\tau}_{\mathrm{p},i}(1+\rho^{2}))\end{split}$ (24a)
$\begin{split}\alpha_{\mathrm{p},i}&<c_{2\rho}\end{split}$ (24b)
$\begin{split}\alpha_{\mathrm{p},i}&<c_{3\rho}\tfrac{\tilde{\tau}_{\mathrm{p},i}(\tilde{\tau}_{\mathrm{p},i}(1+\rho^{2})+2\rho)+1}{4\tilde{\tau}_{\mathrm{p},i}\rho(1+\rho^{2})+2\rho^{2}+5}\end{split}$ (24c)
$\begin{split}\alpha_{\mathrm{p},i}&<c_{4\rho},\end{split}$ (24d)
where $\alpha_{\mathrm{p},i}=\tfrac{d_{\mathrm{p},i}}{\omega_{0}}|v|_{\mathrm{max}}^{2}\sum_{j\neq i}^{n}b_{ij}$, $\tilde{\tau}_{\mathrm{p},i}=\tau_{\mathrm{p},i}\omega_{0}$, and $c_{l\rho}$ for $l=\{1,2,3,4\}$ are quantities depending on $\rho$ as defined in Table 1. For the reactive power-voltage droop control, we require
$\begin{split}\alpha_{\mathrm{q},i}&<c_{5,\rho}(5+10\rho\tilde{\tau}_{\mathrm{q},i})\end{split}$ (25a)
$\begin{split}\alpha_{\mathrm{q},i}&<c_{6,\rho}(2\rho+\tilde{\tau}_{\mathrm{q},i})\end{split}$ (25b)
$\begin{split}\alpha_{\mathrm{q},i}&<c_{7,\rho}\end{split}$ (25c)
$\begin{split}0&<\alpha_{\mathrm{q},i}^{2}c_{8\rho}-\alpha_{\mathrm{q},i}(5\tilde{\tau}_{\mathrm{q},i}(1+\rho^{2})c_{9,\rho}+c_{10,\rho})\\ &\quad\quad\quad+c_{11,\rho}(\tilde{\tau}_{\mathrm{q},i}^{2}(1+\rho^{2})+2\rho\tilde{\tau}_{\mathrm{q},i}+1)\end{split}$ (25d)
$\begin{split}\alpha_{\mathrm{q},i}&<c_{12,\rho}\tilde{\tau}_{\mathrm{q},i}\end{split}$ (25e)
$\begin{split}0&<\alpha_{\mathrm{q},i}^{2}c_{13,\rho}+\alpha_{\mathrm{q},i}\tilde{\tau}_{\mathrm{q},i}c_{14,\rho}+2\tilde{\tau}_{\mathrm{q},i}^{2}\rho\end{split}$ (25f)
where $\alpha_{\mathrm{q},i}=\tfrac{d_{\mathrm{q},i}}{|v|_{0,i}}\sum_{j\neq i}^{n}b_{ij}$, $\tilde{\tau}_{\mathrm{q},i}=\tau_{\mathrm{q},i}\omega_{0}$, and $c_{l\rho}$ for $l=\{5,...,14\}$ are quantities depending on $\rho$ as defined in Table 1.
Remark 3 (Plug-and-Play Operation).
The conditions in 24 and 25 are designed for the scenario where all $n$ converter nodes are populated, ensuring system-level stability during disconnection and re-connection of VSCs. Specifically, removing any VSC decreases the self susceptance $\sum_{j\neq i}^{n}b_{ij}$ for each connected VSC $i$ after Kron reduction and reindexing, thereby guaranteeing that 24 and 25 remain locally satisfied.
[…]
Figure 7: Structural overview of the proof of Theorem 2.
4.1 Main Result
The conditions 24 and 25 are derived in Section 4.2, which proves our main result stated in the theorem below:
Theorem 2 (Internal Feedback Stability of the Closed-Loop System).
Consider the device and network dynamics $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$ in 23, where the network model is given by 17. Let the conditions in 24 and 25 hold for each VSC $i\in\{1,...,n\}$. Then, the closed-loop system $\mathcal{D}_{0}\#\mathcal{N}_{0}$ in Fig. 4 is internally feedback stable, i.e., $\mathcal{D}_{0}\#\mathcal{N}_{0}\in\mathcal{RH}_{\infty}^{4n\times 4n}$.
Corollary 1 (Simplified Network Models).
For the quasi-stationary network dynamics $\mathcal{N}_{0}(s)$ with the network model in 18, where $s=0$ and $|v|_{0,i}\neq|v|_{0,j}$, the stability result of Theorem 2 holds under the relaxed algebraic conditions [^1] These conditions structurally align with the conditions derived in [9].
$\begin{split}\alpha_{\mathrm{q},i}<\tfrac{5(1+\rho^{2})}{4}=c_{7,\rho}\quad\text{and}\quad 0<\tau_{\mathrm{p},i}\end{split}$ (26)
for each VSC $i\in\{1,...,n\}$. Further, for the zero-power-flow network dynamics $\mathcal{N}_{0}(s)$ with the network model in 19, where $s=0$ and $|v|_{0,i}=|v|_{0,j}$, internal feedback stability is guaranteed without imposing any additional VSC conditions.
The proof of Corollary 1 is provided in Appendix 9 and follows a similar reasoning as the proof of Theorem 2 in Section 4.2, but applies the simplified network models in 18 and 19, respectively. It becomes apparent how the conditions in 24 and 25 are a subset of the conditions in 26.
4.2 Proof of Theorem 2
A structural overview of the proof of Theorem 2 is presented in the flowchart in Fig. 7 and consists of four main steps:
I. Coordinate Transformation
To preserve the symmetry of the network, we use the angle and normalized voltage derivatives, defined as $\Delta\omega_{i}(s)\coloneqq\Delta\delta_{i}(s)s$ and $\Delta|\tilde{v}|_{\mathrm{n},i}(s)\coloneqq\Delta|v|_{\mathrm{n},i}(s)s$, as interconnection signals. We then analyze the stability of the closed-loop system as in Fig. 8(a), where
$\begin{split}\mathcal{D}(s)&\coloneqq D(s)\cdot\mathrm{diag}(1,\tfrac{s}{|v|_{0,1}},\dots,1,\tfrac{s}{|v|_{0,n}})\\ \mathcal{N}(s)&\coloneqq N(s)\cdot\mathrm{diag}(\tfrac{1}{s},\tfrac{1}{s},\dots,\tfrac{1}{s},\tfrac{1}{s}).\end{split}$ (27)
II. Loop Shifting
Given that $\mathcal{D}(s)$ and $\mathcal{N}(s)$ in Fig. 8(a) do not satisfy the passivity conditions in Theorem 1 (i.e., it can be shown that $\mathcal{N}(s)$ is not passive), we resort to dynamic loop-shifting techniques as presented in Section 2.3. More specifically, we consider the loop-shifted system $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$ in Fig. 8(b) with the subsystems $\mathcal{D}^{\prime}(s)$ and $\mathcal{N}^{\prime}(s)$ given as
$\begin{split}\mathcal{D}^{\prime}(s)&\coloneqq\mathcal{D}(s)(I-\Gamma(s)\mathcal{D}(s))^{-1}\\ \mathcal{N}^{\prime}(s)&\coloneqq\mathcal{N}(s)+\Gamma(s).\end{split}$ (28)
We choose $\Gamma(s)=\mathrm{diag}(\Gamma_{1}(s),...,\Gamma_{n}(s))$ as a block-diagonal semi-stable transfer matrix with the $2\times 2$ matrix elements
$\Gamma_{i}(s)=\begin{bmatrix}\Gamma_{i}^{\mathrm{p}}(s)&0\\ 0&\Gamma_{i}^{\mathrm{q}}(s)\end{bmatrix},$ (29)
which are selected to ensure passivity of $\mathcal{N}^{\prime}(s)$. Specifically,
$\begin{split}\Gamma_{i}^{\mathrm{p}}(s)&=\tfrac{1}{s}\left(\tfrac{\gamma_{1,i}^{\mathrm{p}}\omega_{0}^{2}s^{2}}{\omega_{0}^{2}+(\rho\omega_{0}+s)^{2}}+\tfrac{\gamma_{2,i}^{\mathrm{p}}\omega_{0}^{2}}{\omega_{0}^{2}+(\rho\omega_{0}+s)^{2}}+\gamma_{3,i}^{\mathrm{p}}\right)\\ \Gamma_{i}^{\mathrm{q}}(s)&=\tfrac{1}{s}\left(\tfrac{\gamma_{1,i}^{\mathrm{q}}\omega_{0}^{2}s^{2}}{\omega_{0}^{2}+(\rho\omega_{0}+s)^{2}}+\tfrac{\gamma_{2,i}^{\mathrm{q}}\omega_{0}^{2}}{\omega_{0}^{2}+(\rho\omega_{0}+s)^{2}}+\gamma_{3,i}^{\mathrm{q}}\right)\end{split}$ (30)
with the parameters
$\begin{split}\gamma_{1,i}^{\mathrm{p}}&=\gamma_{1,i}^{\mathrm{q}}=\tfrac{2|v|_{\mathrm{max}}^{2}}{\omega_{0}^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}\\ \gamma_{2,i}^{\mathrm{p}}&=\gamma_{2,i}^{\mathrm{q}}=-3|v|_{\mathrm{max}}^{2}\textstyle\sum_{j\neq i}^{n}b_{ij}\\ \gamma_{3,i}^{\mathrm{p}}&=3\tfrac{1}{1+\rho^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}|v|_{\mathrm{max}}^{2}\\ \gamma_{3,i}^{\mathrm{q}}&=\tfrac{1}{1+\rho^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}\left(3|v|_{\mathrm{max}}^{2}+0.8\right),\end{split}$ (31)
whose structure loosely resembles the network model in 17.
III. Passivity Checks & Theorem 1
For $\Gamma(s)$ selected as in 29, 30 and 31, we can show that $\mathcal{N}^{\prime}(s)$ is passive [^2] Notice that we can also show passivity of $\mathcal{N}^{\prime}(s)$ for the lossless case $\rho=0$ by evaluating 29, 30 and 31 at $\rho=0$ accordingly., i.e., it satisfies the conditions (i) to (iii) in Definition 2:
[…]
(ii) Positive semi-definiteness: We can express the Hermitian matrix $\mathcal{S}_{\mathcal{N}^{\prime}}(\mathrm{j}\omega)\coloneqq\mathcal{N}^{\prime}(\mathrm{j}\omega)+\mathcal{N}^{\prime\star}(\mathrm{j}\omega)$ as
$\begin{split}\mathcal{S}_{\mathcal{N}^{\prime}}(\mathrm{j}\omega)=\begin{bmatrix}\mathcal{S}_{\mathcal{N}^{\prime},11}(\mathrm{j}\omega)&\dots&\mathcal{S}_{\mathcal{N}^{\prime},1n}(\mathrm{j}\omega)\\ \vdots&\ddots&\vdots\\ \mathcal{S}_{\mathcal{N}^{\prime},n1}(\mathrm{j}\omega)&\dots&\mathcal{S}_{\mathcal{N}^{\prime},nn}(\mathrm{j}\omega)\end{bmatrix},\end{split}$ (32)
where each $\mathcal{S}_{\mathcal{N}^{\prime},ij}$ represents a $2\times 2$ transfer matrix block. The diagonal and off-diagonal elements are given by
$\begin{split}\mathcal{S}_{\mathcal{N}^{\prime},ii}(\mathrm{j}\omega)&=h_{\rho}(\omega)\left(\textstyle\sum_{j\neq i}^{n}|v|_{0,i}^{2}b_{ij}\begin{bmatrix}-1&-\mathrm{j}\omega/\omega_{0}\\ \mathrm{j}\omega/\omega_{0}&-1\end{bmatrix}+\right.\\ &\quad\quad\quad\quad\quad\,\,\,\left.\begin{bmatrix}\omega^{2}\gamma_{1,i}^{\mathrm{p}}-\gamma_{2,i}^{\mathrm{p}}&0\\ 0&\omega^{2}\gamma_{1,i}^{\mathrm{q}}-\gamma_{2,i}^{\mathrm{q}}\end{bmatrix}\right)\\ \mathcal{S}_{\mathcal{N}^{\prime},ij}(\mathrm{j}\omega)&=h_{\rho}(\omega)|v|_{0,i}|v|_{0,j}b_{ij}\begin{bmatrix}1&\mathrm{j}\omega/\omega_{0}\\ -\mathrm{j}\omega/\omega_{0}&1\end{bmatrix}.\end{split}$ (33)
Here, $h_{\rho}(\omega)$ is a positive function $\forall\omega\geq 0$ defined as
$h_{\rho}(\omega)={4\rho\omega_{0}^{3}}/({(\omega_{0}^{2}+\rho^{2}\omega_{0}^{2}-\omega^{2})^{2}+4\rho^{2}\omega_{0}^{2}\omega^{2}})\geq 0.$ (34)
We can observe that 32 is a Hermitian diagonally dominant matrix with real non-negative diagonal entries, i.e., the magnitude of the diagonal entry in a row is greater or equal to the sum of the magnitudes of the off-diagonal entries in that row. Namely, for the odd rows and $\forall\omega\geq 0$, we get
$\begin{split}|\textstyle\sum_{j\neq i}^{n}-|v|_{0,i}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{p}}-\gamma_{2,i}^{\mathrm{p}}|&\geq\\ |\textstyle\sum_{j\neq i}^{n}|v|_{0,i}^{2}b_{ij}\tfrac{-\mathrm{j}\omega}{\omega_{0}}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}&|v|_{0,j}b_{ij}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}|v|_{0,j}b_{ij}\tfrac{\mathrm{j}\omega}{\omega_{0}}|\\ \Leftrightarrow|\textstyle\sum_{j\neq i}^{n}-|v|_{\mathrm{max}}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{p}}-\gamma_{2,i}^{\mathrm{p}}|&\geq\\ 2\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}&b_{ij}\tfrac{\omega}{\omega_{0}}+\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}\hskip 42.67912pt\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(2+\tfrac{2\omega^{2}}{\omega_{0}^{2}})&\geq\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(1+\tfrac{2\omega}{\omega_{0}}),\end{split}$
and similarly, for each even row and $\forall\omega\geq 0$, we have
$\begin{split}|\textstyle\sum_{j\neq i}^{n}-|v|_{0,i}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{q}}-\gamma_{2,i}^{\mathrm{q}}|&\geq\\ |\textstyle\sum_{j\neq i}^{n}|v|_{0,i}^{2}b_{ij}\tfrac{\mathrm{j}\omega}{\omega_{0}}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}&|v|_{0,j}b_{ij}\tfrac{-\mathrm{j}\omega}{\omega_{0}}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}|v|_{0,j}b_{ij}|\\ \Leftrightarrow|\textstyle\sum_{j\neq i}^{n}-|v|_{\mathrm{max}}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{q}}-\gamma_{2,i}^{\mathrm{q}}|&\geq\\ 2\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}&b_{ij}\tfrac{\omega}{\omega_{0}}+\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(2+\tfrac{2\omega^{2}}{\omega_{0}^{2}})&\geq\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(1+\tfrac{2\omega}{\omega_{0}}).\end{split}$
By Lemma 1, we can conclude $\mathcal{N}^{\prime}(\mathrm{j}\omega)+\mathcal{N}^{\prime\star}(\mathrm{j}\omega)\succeq 0$.
$\mathcal{D}(s)$$\mathcal{N}(s)$--$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|\tilde{v}|_{\mathrm{n}}\end{bmatrix}$$\begin{bmatrix}\Delta p\\ \Delta q\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|\tilde{v}|_{\mathrm{nd}}\end{bmatrix}$device dynamicsnetwork dynamics--$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$
(a) Original feedback system.
$\mathcal{D}(s)$$\mathcal{D}^{\prime}$$\mathcal{N}(s)$$\mathcal{N}^{\prime}$--$\Gamma(s)$$\Gamma(s)$$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|\tilde{v}|_{\mathrm{n}}\end{bmatrix}$device dynamicsnetwork dynamics$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|\tilde{v}|_{\mathrm{nd}}\end{bmatrix}$
(b) Loop-shifting with $\Gamma$.
Figure 8: Closed-loop feedback interconnection for stability analysis.
[…]
$\begin{split}\mathcal{R}_{\mathrm{j}0,ii}^{\mathcal{N}^{\prime}}&=\textstyle\sum_{j\neq i}^{n}b_{ij}\begin{bmatrix}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}&0\\ 0&\tfrac{2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\end{bmatrix}+\\ &\quad\quad\quad\quad\quad\quad\quad\quad\,\,\,\begin{bmatrix}\tfrac{\gamma_{2,i}^{\mathrm{p}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{p}}&0\\ 0&\tfrac{\gamma_{2,i}^{\mathrm{q}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{q}}\end{bmatrix}\\ \mathcal{R}_{\mathrm{j}0,ij}^{\mathcal{N}^{\prime}}&=-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}.\end{split}$ (35)
Again, 35 is a Hermitian diagonally dominant matrix with real non-negative diagonal entries. For the odd rows, we get
$\begin{split}|\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}+\tfrac{\gamma_{2,i}^{\mathrm{p}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{p}}|&\geq\textstyle\sum_{j\neq i}^{n}|-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}&\geq\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}.\end{split}$
For the even rows we get (with $|v|_{\mathrm{max}}=1.1$ and $|v|_{\mathrm{min}}=0.9$):
$\begin{split}|\textstyle\sum_{j\neq i}^{n}\tfrac{(2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j})b_{ij}}{1+\rho^{2}}+\tfrac{\gamma_{2,i}^{\mathrm{q}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{q}}|&\geq\textstyle\sum_{j\neq i}^{n}|b_{ij}\tfrac{-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\\ \Leftrightarrow|\textstyle\sum_{j\neq i}^{n}\tfrac{(2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}+0.8)b_{ij}}{1+\rho^{2}}|&\geq\textstyle\sum_{j\neq i}^{n}|b_{ij}\tfrac{-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{2|v|_{\mathrm{min}}^{2}-|v|_{\mathrm{max}}^{2}+0.8}{1+\rho^{2}}&\geq\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{\mathrm{max}^{2}}}{1+\rho^{2}}\end{split}$
By Lemma 1, we conclude $\mathcal{R}_{\mathrm{j}0}^{\mathcal{N}^{\prime}}\succeq 0$.
Next, we derive the decentralized stability conditions in 24 and 25 under which $\mathcal{D}^{\prime}(s)$ is strictly passive. We consider $\mathcal{D}^{\prime}(s)=\text{diag}(\mathcal{D}_{1}^{\prime}(s),\dots,\mathcal{D}_{n}^{\prime}(s))$ with the matrix elements
$\begin{split}\mathcal{D}_{i}^{\prime}(s)=\mathcal{D}_{i}(s)(I-\Gamma_{i}(s)\mathcal{D}_{i}(s))^{-1}=\begin{bmatrix}\mathcal{D}_{\mathrm{p},i}^{\prime}(s)&0\\ 0&\mathcal{D}^{\prime}_{\mathrm{q},i}(s)\end{bmatrix},\end{split}$ (36)
where the diagonal transfer function elements are given as
$\begin{split}\mathcal{D}_{\mathrm{p},i}^{\prime}(s)&=\tfrac{d_{\mathrm{p},i}}{\tau_{\mathrm{p},i}s+1-\Gamma_{i}^{\mathrm{p}}(s)d_{\mathrm{p},i}}\end{split}$ (37a)
$\begin{split}\mathcal{D}_{\mathrm{q},i}^{\prime}(s)&=\tfrac{d_{\mathrm{q},i}s/|v|_{0,i}}{\tau_{\mathrm{q},i}s+1-\Gamma_{i}^{\mathrm{q}}(s)d_{\mathrm{q},i}s/|v|_{0,i}},\end{split}$ (37b)
for each of which we check the strict passivity conditions (i) and (ii) in Definition 3 independently. We start with $\mathcal{D}_{\mathrm{p},i}^{\prime}(s)$:
(i) Poles: To show that the poles of all elements of $\mathcal{D}_{\mathrm{p},i}^{\prime}(s)$ are in $\text{Re}(s)<0$, we start by inserting the expression for $\Gamma_{i}^{\mathrm{p}}(s)$ in 30 into 37a and obtain
$\mathcal{D}_{\mathrm{p},i}^{\prime}(s)=d_{\mathrm{p},i}\tfrac{s^{2}b_{2,i}+sb_{1,i}+b_{0,i}}{s^{3}a_{3,i}+s^{2}a_{2,i}+sa_{1,i}+a_{0,i}},$ (38)
where the transfer function coefficients are given by
$\begin{split}a_{0,i}&=\omega_{0}^{2}(\rho^{4}+2\rho^{2}-6\rho\alpha_{\mathrm{p},i}+1)\\ a_{1,i}&=\omega_{0}(\tilde{\tau}_{\mathrm{p},i}\rho^{4}+2\rho^{3}+2\rho^{2}\tilde{\tau}_{\mathrm{p},i}\\ &-2\rho^{2}\alpha_{\mathrm{p},i}+2\rho-5\alpha_{\mathrm{p},i}+\tilde{\tau}_{\mathrm{p},i})\\ a_{2,i}&=(1+\rho^{2})(2\tilde{\tau}_{\mathrm{p},i}\rho+1)\\ a_{3,i}&=\tau_{\mathrm{p},i}(1+\rho^{2})\end{split}\begin{split}b_{0,i}&=\omega_{0}^{2}(1+\rho^{2})^{2}\\ b_{1,i}&=2\rho\omega_{0}(1+\rho^{2})\\ b_{2,i}&=1+\rho^{2}.\end{split}$ (39)
By checking the Hurwitz criterion for a 3rd order polynomial, we can conclude that $a_{3,i}>0$ and $a_{2,i}>0$ are always satisfied, while $a_{1,i}>0$ is ensured by condition 24a, $a_{0,i}>0$ by condition 24b, and $a_{2,i}a_{1,i}>a_{0,i}a_{3,i}$ by condition 24c.
(ii) Positive definiteness: To ensure $\mathcal{D}_{\mathrm{p},i}^{\prime}(\mathrm{j}\omega)+\mathcal{D}^{\prime\star}_{\mathrm{p},i}(\mathrm{j}\omega)\succ 0$, we require the additional condition 24d.
Following the same reasoning, we derive the decentralized stability conditions [^3] The requirement $\tau_{\mathrm{q},i}\neq 0$ in 25 ensures properness of $\mathcal{D}_{\mathrm{q},i}^{\prime}(s)$. for the reactive power-voltage droop control presented in 25 by checking strict passivity of $\mathcal{D}_{\mathrm{q},i}^{\prime}(s)$.
Finally, since $\mathcal{N}^{\prime}(s)$ is passive and $\mathcal{D}^{\prime}(s)$ is strictly passive, we can apply Theorem 1 and conclude $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}\in\mathcal{RH}_{\infty}^{4n\times 4n}$.
IV. Final Value Theorem & Lemma 2
The internal feedback stability of the original system $\mathcal{D}_{0}\#\mathcal{N}_{0}$ follows directly from the internal feedback stability of the loop-shifted system $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$, as stated in the following lemma.
Lemma 4.
Consider the device and network dynamics $\mathcal{D}^{\prime}(s)$ and $\mathcal{N}^{\prime}(s)$ in 28, as well as $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$ in 23 with the network model 17. Then, internal feedback stability of $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$ implies internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$.
Proof.
We first conclude that $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}\in\mathcal{RH}_{\infty}^{4n\times 4n}$ implies $(I+\mathcal{D}^{\prime}\mathcal{N}^{\prime})^{-1}\mathcal{D}^{\prime}=(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}\in\mathcal{RH}_{\infty}^{2n\times 2n}$. Further, since
$\begin{split}\mathcal{D}(s)&=\mathcal{D}_{0}(s)\cdot\text{diag}(1,\tfrac{s}{|v|_{0,1}},\dots,1,\tfrac{s}{|v|_{0,n}})\\ \mathcal{N}(s)&=\mathcal{N}_{0}(s)\cdot\text{diag}(1,\tfrac{|v|_{0,1}}{s},\dots,1,\tfrac{|v|_{0,n}}{s}),\end{split}$ (40)
$(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}=\text{diag}(1,\tfrac{s}{|v|_{0,1}},\dots,1,\tfrac{s}{|v|_{0,n}})(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}$, and we can thus conclude stability of $(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}$ if the step response of the voltage derivatives converges to zero, i.e.,
$\underset{s\rightarrow 0}{\text{lim}}\,\,s\frac{1}{s}(I+\mathcal{D}(s)\mathcal{N}(s))^{-1}\mathcal{D}(s)=I\otimes\begin{bmatrix}\star&0\\ 0&0\end{bmatrix}.$ (41)
To show that 41 holds, we first rewrite the transfer matrix $(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}=(I+\mathcal{D}{N}\tfrac{1}{s})^{-1}\mathcal{D}$, and then use the row permutation matrix $\mathcal{P}\in\mathbb{R}^{2n\times 2n}$ with elements
$\mathcal{P}_{ij}=\begin{cases}1\quad\quad i=k,\,j=2k-1,\,1\leq k\leq n\\ 1\quad\quad i=k+n,\,j=2k,\,1\leq k\leq n\\ 0\quad\quad\text{else},\end{cases}$ (42)
to study the decoupled frequency and voltage dynamics, i.e., $\mathcal{P}(I+\mathcal{D}{N}\tfrac{1}{s})^{-1}\mathcal{D}\mathcal{P}^{-1}=(I+\mathcal{D}^{\mathcal{P}}{N}^{\mathcal{P}}\tfrac{1}{s})^{-1}\mathcal{D}^{\mathcal{P}}$. The row-permutated matrices $\mathcal{D}^{\mathcal{P}}(s)$ and $N^{\mathcal{P}}(s)$ are given as
$\mathcal{D}^{\mathcal{P}}(s)=\begin{bmatrix}\mathcal{D}_{\mathrm{p}}^{\mathcal{P}}(s)&0_{n\times n}\\ 0_{n\times n}&\mathcal{D}_{\mathrm{q}}^{\mathcal{P}}(s)\end{bmatrix},\,\,N^{\mathcal{P}}(s)=\begin{bmatrix}N_{1}^{\mathcal{P}}(s)&N_{2}^{\mathcal{P}}(s)\\ -N_{2}^{\mathcal{P}}(s)&N_{3}^{\mathcal{P}}(s)\end{bmatrix}$ (43)
where $\mathcal{D}_{\mathrm{p}}^{\mathcal{P}}(s)=\text{diag}(\mathcal{D}_{\mathrm{p},1}(s),\dots,\mathcal{D}_{\mathrm{p},n}(s))$ and $\mathcal{D}_{\mathrm{q}}^{\mathcal{P}}(s)=\text{diag}(\mathcal{D}_{\mathrm{q},1}(s),\dots,\mathcal{D}_{\mathrm{q},n}(s))$ with elements
$\mathcal{D}_{\mathrm{p},i}(s)=\tfrac{d_{\mathrm{p},i}}{\tau_{\mathrm{p},i}s+1}\quad\text{and}\quad\mathcal{D}_{\mathrm{q},i}(s)=\tfrac{d_{\mathrm{q},i}s}{\tau_{\mathrm{q},i}s+1}\tfrac{1}{|v|_{0,i}}.$ (44)
Moreover, we have
$N_{1,ii}^{\mathcal{P}}(s)$ $=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}-\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}$
$N_{1,ij}^{\mathcal{P}}(s)$ $=-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}$
$\begin{split}N_{2,ii}^{\mathcal{P}}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\tfrac{s}{\omega_{0}}\\ N_{2,ij}^{\mathcal{P}}(s)&=-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\tfrac{s}{\omega_{0}}\end{split}$ (45)
$N_{3,ii}^{\mathcal{P}}(s)$ $=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}+\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}$
$N_{3,ij}^{\mathcal{P}}(s)$ $=-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}.$
Recalling that $(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}$ is stable, and thus its permuted version, we can apply the Final Value Theorem (FVT):
$\begin{split}&\,\,\underset{s\rightarrow 0}{\text{lim}}\,\,(I+\mathcal{D}^{\mathcal{P}}(s){N}^{\mathcal{P}}(s)\tfrac{1}{s})^{-1}\mathcal{D}^{\mathcal{P}}(s)\\ =\,&\,\,\underset{s\rightarrow 0}{\text{lim}}\,\,(sI+\mathcal{D}^{\mathcal{P}}(s){N}^{\mathcal{P}}(s))^{-1}\mathcal{D}^{\mathcal{P}}(s)s\\ =\,&\,\,\underset{s\rightarrow 0}{\text{lim}}\,\left[\begin{array}[]{c}(sI+\mathcal{D}^{\mathcal{P}}_{\mathrm{p}}(s){N}^{\mathcal{P}}_{1}(s))^{-1}\mathcal{D}^{\mathcal{P}}_{\mathrm{p}}(s)s\\ 0_{n\times n}\end{array}\right.\\ \quad&\hskip 71.13188pt\left.\begin{array}[]{c}0_{n\times n}\\ (sI+\mathcal{D}^{\mathcal{P}}_{\mathrm{q}}(s){N}^{\mathcal{P}}_{3}(s))^{-1}\mathcal{D}^{\mathcal{P}}_{\mathrm{q}}(s)s\end{array}\right]\\ =\,&\begin{bmatrix}\star_{n\times n}&0_{n\times n}\\ 0_{n\times n}&0_{n\times n}\end{bmatrix},\end{split}$ (46)
where for the second equality we have used $N_{2}^{\mathcal{P}}(0)=0_{n\times n}$. The last equality follows from the fact that $N_{1}^{\mathcal{P}}(0)$ is a Laplacian matrix with zero eigenvalue [22], while $N_{3}^{\mathcal{P}}(0)$ is a regular matrix which does not have a zero eigenvalue. Given 46, 41 follows, and thus $(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}\in\mathcal{RH}_{\infty}^{2n\times 2n}$.
Finally, since $\mathcal{D}_{0}(s)$ does not include a RHP pole or zero, there are no RHP pole-zero cancellations between $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$. By Lemma 2, we conclude $\mathcal{D}_{0}\#\mathcal{N}_{0}\in\mathcal{RH}_{\infty}^{4n\times 4n}$. ∎
(a) Stability conditions in 24.
(b) Stability conditions in 25.
Figure 9: 2D plot of the stability conditions for the three-node system with $\rho=0.05$. The stars indicate the feasibility of the GFM 3 control parameters.
5 Numerical Case Studies
5.1 Case Study I: Linearized & reduced models
We numerically validate Theorem 2 and Corollary 1 with MATLAB/Simulink using linearized small-signal models of a three-node system with three GFM VSCs. We consider a network with uniform resistance-inductance ratio $\rho=0.05$, a maximum steady-state voltage magnitude $|v|_{\mathrm{max}}=1.1$ pu, and identical self susceptances $\sum_{i\neq j}^{3}b_{ij}=5$ pu of all VSC nodes $i\in\{1,2,3\}$. The simulations follow the ideal (i.e., linearized and reduced) block diagram dynamics in Fig. 4. The nodes GFM 1 and GFM 2 employ fixed controllers that always satisfy the conditions in 24 and 25, while for node GFM 3 we are exploring varying control parameters.
We begin by modeling the network dynamics $\mathcal{N}_{0}(s)$ using Network-Simplification Level 2 in 19, without imposing additional conditions on the controller of GFM 3 (see “no cond.” in Fig. 9). A small-signal load disturbance at node 1 reveals that the closed-loop system remains stable (Fig. 10), confirming the validity of Corollary 1. Likewise, when modeling $\mathcal{N}_{0}(s)$ with Network-Simplification Level 1 in 18 and ensuring that the controller of GFM 3 satisfies the stability conditions in 26 (see “cond. L1” in Fig. 9), we again observe closed-loop stability (Fig. 11). This further confirms Corollary 1. Our main result, Theorem 2, is validated in Fig. 12, where we model $\mathcal{N}_{0}(s)$ dynamically as in 17, and equip GFM 3 with a controller that meets the stability conditions in 24 and 25 (see “cond. DYN” in Fig. 9). Stability is immediately evident. In contrast, using a GFM 3 controller that satisfies only 26 (“cond. L1”) or no conditions at all (“no cond.”) leads to instability in the dynamic network model. This underscores the importance of accurate network modeling in control design, which is overlooked in the overly optimistic stability assessment of [7, 9, 10].
5.2 Case Study II: Nonlinear circuit model
For the same three-node system as in Section 5.1, we now use Simscape Electrical in MATLAB/Simulink to perform a detailed electromagnetic transients (EMT) simulation based on a nonlinear circuit model of the three-phase network. Additionally, for each GFM VSC, we incorporate the full nonlinear converter models as shown in Fig. 3. Compared to Case Study I, this increases complexity in two key aspects: first, by considering nonlinear models, and second, by accounting for the full network dynamics without the simplifying assumptions introduced in Section 3.1. The simulation results in Fig. 13 demonstrate that a GFM 3 controller satisfying the decentralized stability conditions in 24 and 25 ensures stability even in the presence of nonlinear network and device models.
Figure 10: Stable system response of the block diagram in Fig. 4 with three GFM devices and the simplified network model 19, where the controller of GFM 3 does not satisfy any of the conditions 24, 25 and 26 (no cond.).
Figure 11: Stable system response of the block diagram in Fig. 4 with three GFM devices and the simplified network model 18, where the GFM 3 controller satisfies the conditions in 26 (cond. L1).
Figure 12: Stable system response of the block diagram in Fig. 4 with three GFM devices and the dynamic network model 17, where all GFM controllers are satisfying the conditions in 24 and 25 (cond. DYN). We also display the unstable dynamics of GFM 3 when the controller is satisfying no conditions (no cond.), and when it is satisfying the conditions in 26 (cond. L1).
Figure 13: Stable system response of the three-node converter system with nonlinear circuit and device models where all GFM controllers are satisfying the conditions in 24 and 25 (cond. DYN). Additionally, we indicate the unstable dynamics of GFM 3 when the controller is satisfying no conditions (no cond.), and when it is satisfying the conditions in 26 (cond. L1).
In contrast, when the GFM 3 controller satisfies only 26 or does not meet any stability conditions, interactions with the network dynamics can lead to instability. Overall, we conclude that our stability conditions in 24 and 25, which explicitly account for network dynamics, remain effective in a nonlinear circuit scenario, provided the system operates near the nominal point where linearization errors are sufficiently small. However, the stability conditions, designed for a simplified static network, may fail when network dynamics are taken into account.
6 Conclusion
We proposed a decentralized small-signal stability certification framework to mitigate the destabilizing effects of network dynamics on grid-forming converters. Using dynamic loop-shifting techniques and passivity theory, we derived parametric stability conditions that can serve as local tuning rules for device-level controllers, eliminating the need for centralized coordination. Our numerical case studies validated our theoretical results even for non-linear circuit and device models.
Future work includes the extension of our dynamic loop-shifting framework to a non-diagonal structure of the matrix blocks $\Gamma_{i}(s)$ with cross-coupling between active and reactive power, thereby eventually reducing conservatism by requesting coupled device controllers. Beyond that, from an application point of view, we envision our stability framework to support the formulation of new grid codes for future power systems.

### v6

@@
 […]
-The transition to future power systems is characterized by a substantial increase in the share of power electronics (PE)-based generation devices.
-This shift introduces significant changes in system dynamics, where the interactions between PE-based generation and the remainder of the power network are not fully understood yet [1].
-In particular, the fast and complex dynamics of PE-based devices can lead to unexpected effects (e.g., overvoltages [2], subsynchronous oscillations [3], etc.), posing considerable challenges for maintaining system stability.
-Consequently, there is a pressing need for a deeper understanding of how PE-based devices interact with one another or with the grid and for the development of robust stability frameworks to ensure stable operation.
+The transition to future power systems is characterized by a substantial increase in the share of power electronics (PE)-based generation devices [1].
+This shift introduces significant changes in system dynamics, where the interactions between PE-based generation and the remainder of the power network are not fully understood yet [2, 3, 4].
+In particular, the faster dynamics of the control loops and filters in PE-based devices, compared to conventional generators, may induce unforeseen phenomena (e.g., overvoltages[3], subsynchronous oscillations[4]), thereby posing significant challenges to system stability.
+Consequently, there is a pressing need for a deeper understanding of how PE-based devices interact with one another or with the grid and for the development of stability frameworks to ensure stable operation.
 One of the key aspects of future power systems involves the integration of PE-based devices with either grid-following or grid-forming control.
@@
 By formulating device-level specifications that are sufficiently passive, we ensure overall system stability in a flexible and scalable way.
-Our contribution improves significantly over prior works on stability certification in GFM converter systems.
+Our contribution improves significantly over prior works on stability certification of GFM converter systems in the small-signal regime.
 Unlike [8, 9], which are limited to single-input single-output (SISO) dynamics, our approach accommodates multiple-input multiple-output (MIMO) dynamics coupling frequency and voltage, thus providing a more comprehensive analysis of the system behavior.
@@
 Moreover, unlike [13], our approach does not require detailed knowledge of the network configuration.
-Finally, our stability conditions are independent of the operating point, provided small bus angle differences, unlike [9], where the the stability conditions depend on specific operating points.
 Collectively, these attributes allow our framework to provide, to the best of our knowledge, the most general, explicitly parametric, and practically applicable small-signal stability certificates for interconnected GFM devices in the literature.
@@
 We use $A=\mathrm{diag}(A_{1},A_{2},...,A_{k}),\,k\in\mathbb{Z}$ to denote the block-diagonal matrix with blocks $A_{1},A_{2},...,A_{k}$.
-The Euclidean norm of a matrix $A$ is defined as $||A||_{F}=\sqrt{\textstyle\sum_{i,j}|a_{ij}|^{2}}$, where $a_{ij}$ is the entry in the $i$th row and $j$th column of $A$.
+The Euclidean norm of a matrix $A$ is defined as $A||_{F}=\sqrt{\textstyle\sum_{i,j}|a_{ij}|^{2}}$, where $a_{ij}$ is the entry in the $i$th row and $j$th column of $A$.
 For a complex matrix $A\in\mathbb{C}^{n\times n}$, we use $A^{\star}$ to denote its conjugate transpose.
@@
 A matrix $A\in\mathbb{C}^{n\times n}$ is called diagonally dominant, if $|a_{ii}|\geq\textstyle\sum_{j\neq i}|a_{ij}|,\,\forall i$.
-To establish our main stability results in Section 4, we require the following lemma which follows from the Gershgorin’s Circle Theorem [13, Thm. 6.1.10]:
+To establish our main stability results in Section 4, we require the following lemma which follows from the Gershgorin’s Circle Theorem [14, Thm. 6.1.10]:
 Lemma 1.
 A Hermitian diagonally dominant matrix with real non-negative diagonal entries is positive semi-definite.
-Further, we review the recent concept of phases of complex matrices based on the matrix’s numerical range [14].
-Namely, the numerical range of a complex matrix $A\in\mathbb{C}^{n\times n}$ is defined as $W(A)=\{x^{\star}Ax:x\in\mathbb{C}^{n},||x||=1\}$.
-If $0\notin W(A)$, then $A$ is said to be a sectorial matrix.
-For a sectorial $A$, there exists a nonsingular matrix $T$ and a diagonal unitary matrix $D$ such that $A=T^{\star}DT$ [14].
-The diagonal elements of $D$ are distributed in an arc on the unit circle with length smaller than $\pi$.
-Then, the phases of $A$, denoted by
+[…]
 $\overline{\phi}(A)=\phi_{1}(A)\geq\dots\geq\phi_{n}(A)=\underline{\phi}(A),$ (1)
@@
 Let $u(t)\in\mathbb{R}^{n}$ be the input and $y(t)\in\mathbb{R}^{n}$ the output signal of a MIMO LTI system.
-The transfer matrix $H(s)\in\mathbb{C}^{n\times n}$ describes the input-output system response in the frequency domain as $y(s)=H(s)u(s)$, where $y(s)$ and $u(s)$ are the Laplace transformations of the output and input, respectively.
+The $n\times n$ transfer matrix $H(s)$ describes the input-output system response in the frequency domain as $y(s)=H(s)u(s)$, where $y(s)$ and $u(s)$ are the Laplace transformations of the output and input, respectively.
 We denote the space of $n$-by-$n$ real rational proper transfer matrices of stable LTI systems by $\mathcal{RH}_{\infty}^{n\times n}$.
@@
 1(a) is internally feedback stable if and only if the following four closed-loop transfer matrices
-$\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}=\underset{\eqqcolon\,H_{1}\#H_{2}(s)}{\underbrace{\begin{bmatrix}(I+H_{1}H_{2})^{-1}H_{1}&-(I+H_{1}H_{2})^{-1}H_{1}H_{2}\\ H_{2}(I+H_{1}H_{2})^{-1}H_{1}&H_{2}(I+H_{1}H_{2})^{-1}\end{bmatrix}}}\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}$ (2)
+$\hskip-7.11317pt\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix}\hskip-3.41432pt=\hskip-3.41432pt\underset{\eqqcolon\,H_{1}\#H_{2}(s)}{\underbrace{\begin{bmatrix}(I\hskip-1.70717pt+\hskip-1.70717ptH_{1}H_{2})^{-1}H_{1}&-(I\hskip-1.70717pt+\hskip-1.70717ptH_{1}H_{2})^{-1}H_{1}H_{2}\\ H_{2}(I\hskip-1.70717pt+\hskip-1.70717ptH_{1}H_{2})^{-1}H_{1}&H_{2}(I\hskip-1.70717pt+\hskip-1.70717ptH_{1}H_{2})^{-1}\end{bmatrix}}}\hskip-4.2679pt\begin{bmatrix}w_{1}\\ w_{2}\end{bmatrix}\hskip-4.55244pt$ (2)
 are stable, compactly referred to as $H_{1}\#H_{2}\in\mathcal{RH}_{\infty}^{2n\times 2n}$.
-If there are no RHP pole-zero cancellations between $H_{1}(s)$ and $H_{2}(s)$, then stability of one closed-loop transfer matrix implies stability of the others [15, Thm 4.7]:
+If there are no RHP pole-zero cancellations between $H_{1}(s)$ and $H_{2}(s)$, then stability of one closed-loop transfer matrix implies stability of the others [16, Thm 4.5]:
 Lemma 2.
@@
 1(a) is internally feedback stable if and only if one of the four closed-loop transfer matrices in 2 is stable.
-$H_{1}(s)$$H_{2}(s)$--$w_{1}$$y_{1}$$y_{2}$$w_{2}$
+$H_{1}(s)$$H_{2}(s)$-$w_{1}$$y_{1}$$y_{2}$$w_{2}$
 (a) Original feedback system.
-[…]
+$H_{1}(s)$$H^{\prime}_{1}$$H_{2}(s)$$H^{\prime}_{2}$-$w_{1}$$y_{1}$$\Gamma(s)$$\Gamma(s)$$y_{2}$$w_{2}$
 (b) Loop shifting with $\Gamma(s)$.
@@
 2.3 Small-Phase & Passivity Theory
-For LTI systems, the property of passivity is equivalent to the property of positive realness [14, 16, 17, 18].
-Definition 2 (Passive Transfer Matrix [14]).
-A $n\times n$ real rational proper transfer matrix $H(s)$ is passive if
+For LTI systems, the property of passivity is equivalent to the property of positive realness [17, 18, 19].
+Definition 2 (Passivity).
+A $n\times n$ real rational proper transfer matrix $H(s)$ is said to be passive if
   1. (i)
@@
   3. (iii)
-    any purely imaginary pole $\mathrm{j}\omega$ of any element of $H(s)$ is a simple pole and $\text{lim}_{s\rightarrow\mathrm{j}\omega}(s-\mathrm{j}\omega)H(s)$ is positive semi-definite Hermitian.
-Definition 3 (Strictly Passive Transfer Matrix [14]).
-A $n\times n$ real rational proper transfer matrix $H(s)$ is strictly passive if
+    any purely imaginary pole $\mathrm{j}\omega$ of any element of $H(s)$ is a simple pole and $\text{lim}_{s\rightarrow\mathrm{j}\omega}(s\hskip-1.99168pt-\hskip-1.99168pt\mathrm{j}\omega)H(s)$ is positive semi-definite Hermitian.
+Definition 3 (Strict Passivity).
+A $n\times n$ real rational proper transfer matrix $H(s)$ is said to be strictly passive if
   1. (i)
@@
   2. (ii)
-    $H(\mathrm{j}\omega)+H^{\star}(\mathrm{j}\omega)\succ 0$ for any $\omega\in[-\infty,\infty]$.
-We now present a generalization of the passivity theorem [17, 18] for the special case of LTI systems as follows:
+    $H(\mathrm{j}\omega)+H^{\star}(\mathrm{j}\omega)\succ 0$ for any $\omega\in(-\infty,\infty)$.
+We now present a variant of the passivity theorem [19] for LTI systems as follows:
 Theorem 1 (Passivity Theorem).
-Consider two LTI systems $H_{1}(s)$ and $H_{2}(s)$ in negative feedback configuration, as shown in Fig.
-1(a).
-The feedback system is internally feedback stable if $H_{1}(s)$ is strictly passive and $H_{2}(s)$ is passive.
-The above theorem is widely known in the SISO case [17], but we could not locate a proof in this general MIMO setting.
-We thus provide a self-contained proof below, relying on the Generalized Small-Phase Theorem [14, Thm 7.1]:
-Lemma 3 (Generalized Small-Phase Theorem).
-Let $H_{1}(s)\in\mathcal{RH}_{\infty}^{n\times n}$ be frequency-wise sectorial and $H_{2}(s)$ be semi-stable frequency-wise semi-sectorial with $\mathrm{j}\Omega$ being the set of poles on the imaginary axis.
+Consider two LTI systems $H_{1}(s)$ and $H_{2}(s)$ in negative feedback configuration, as in Fig. 1(a). The feedback system is internally feedback stable if
+  • $H_{1}(s)$ is strictly passive,
+  • $H_{2}(s)$ is passive, and
+  • $\bar{\sigma}(H_{1}(\mathrm{j}\infty))\bar{\sigma}(H_{2}(\mathrm{j}\infty))<1$.
+Several variants of the above theorem exist, often for the SISO case, with subtle distinctions arising particularly in the special case $\omega=\infty$.
+To avoid ambiguity, we provide a self-contained proof below, relying on the Mixed Gain–Phase Theorem with Cutoff Frequency [20].
+Lemma 3 (Mixed Gain-Phase Theorem with Cutoff Frequency).
+Let $\omega_{\mathrm{c}}\in(0,\infty)$, $H_{1}(s)\in\mathcal{RH}_{\infty}^{n\times n}$ be frequency-wise sectorial, and $H_{2}(s)$ be semi-stable frequency-wise semi-sectorial over $(-\omega_{\mathrm{c}},\omega_{\mathrm{c}})$ with $\mathrm{j}\Omega$ being the set of poles on the imaginary axis satisfying $\text{max}_{\omega\in\Omega}|\omega|<\omega_{\mathrm{c}}$.
 Then, $H_{1}\#H_{2}$ in Fig.
 1(a) is internally feedback stable if
-$\begin{split}&\underset{\omega\in[0,\infty]\backslash\Omega}{\mathrm{sup}}\left[\overline{\phi}(H_{1}(\mathrm{j}\omega))+\overline{\phi}(H_{2}(\mathrm{j}\omega))\right]<\pi\quad\text{and}\\ &\underset{\omega\in[0,\infty]\backslash\Omega}{\mathrm{inf}}\left[\underline{\phi}(H_{1}(\mathrm{j}\omega))+\underline{\phi}(H_{2}(\mathrm{j}\omega))\right]>-\pi.\end{split}$ (3)
+  1. (i)
+    for each $\omega\in[0,\omega_{\mathrm{c}})\backslash\Omega$, it holds
+    $\begin{split}&\overline{\phi}(H_{1}(\mathrm{j}\omega))+\overline{\phi}(H_{2}(\mathrm{j}\omega))<\pi\quad\text{and}\\ &\underline{\phi}(H_{1}(\mathrm{j}\omega))+\underline{\phi}(H_{2}(\mathrm{j}\omega))>-\pi,\end{split}$ (3)
+  2. (ii)
+    and for each $\omega\in[\omega_{\mathrm{c}},\infty]$, it holds
+    $\bar{\sigma}(H_{1}(\mathrm{j\omega}))\bar{\sigma}(H_{2}(\mathrm{j\omega}))<1.$ (4)
 Proof of Theorem 1.
-We prove Theorem 1 by applying Lemma 3 as follows: Since $H_{1}(s)$ is strictly passive, it follows that $H_{1}(s)\in\mathcal{RH}_{\infty}^{n\times n}$ is frequency-wise sectorial with phases $\left[\underline{\phi}(H_{1}(\mathrm{j}\omega)),\overline{\phi}(H_{1}(\mathrm{j}\omega))\right]\subset(-\tfrac{\pi}{2},\tfrac{\pi}{2}),\forall\omega\in[0,\infty]$ (see [14] for a detailed derivation).
-Likewise, since $H_{2}(s)$ is passive, it follows that $H_{2}(s)$ is semi-stable with poles at $\mathrm{j}\Omega$ on the imaginary axis and frequency-wise semi-sectorial with phases $\left[\underline{\phi}(H_{2}(\mathrm{j}\omega)),\overline{\phi}(H_{2}(\mathrm{j}\omega))\right]\subset[-\tfrac{\pi}{2},\tfrac{\pi}{2}],\forall\omega\in[0,\infty]\backslash\Omega$ (see [14] for a detailed derivation).
-Therefore, the conditions in 3 are satisfied which implies $H_{1}\#H_{2}\in\mathcal{RH}_{\infty}^{2n\times 2n}$.
+We establish Theorem 1 by invoking Lemma 3 with $\omega_{\mathrm{c}}=\infty$.
+Since Lemma 3 holds for arbitrarily large $\omega_{\mathrm{c}}$, continuity ensures its validity also at $\omega_{\mathrm{c}}=\infty$.
+Accordingly, we proceed as follows.
+Because $H_{1}(s)$ is strictly passive, it follows that $H_{1}(s)\in\mathcal{RH}_{\infty}^{n\times n}$ is frequency-wise sectorial with phases $\left[\underline{\phi}(H_{1}(\mathrm{j}\omega)),\overline{\phi}(H_{1}(\mathrm{j}\omega))\right]\subset(-\tfrac{\pi}{2},\tfrac{\pi}{2}),\forall\omega\in[0,\infty)$.
+Likewise, since $H_{2}(s)$ is passive, it follows that $H_{2}(s)$ is semi-stable with poles at $\mathrm{j}\Omega$ on the imaginary axis and frequency-wise semi-sectorial with phases $\left[\underline{\phi}(H_{2}(\mathrm{j}\omega)),\overline{\phi}(H_{2}(\mathrm{j}\omega))\right]\subset[-\tfrac{\pi}{2},\tfrac{\pi}{2}],\forall\omega\in[0,\infty]\backslash\Omega$.
+Hence, the conditions in 3 in Lemma 3 are satisfied.
+For $\omega=\infty$, it becomes immediate how 4 in Lemma 3 is satisfied.
+This implies $H_{1}\#H_{2}\in\mathcal{RH}_{\infty}^{2n\times 2n}$.
 ∎
@@
 1(a) are not satisfying the passivity conditions in Theorem 1, the two subsystems ${H}^{\prime}_{2}(s)$ and ${H}^{\prime}_{1}(s)$ in Fig.
-1(b) might do so for a suitable $\Gamma(s)$, thus resulting in a stable feedback interconnection ${H}^{\prime}_{1}\#{H}^{\prime}_{2}$ (for details see literature on frequency-dependent passivity indices, e.g., [17, 18]).
+1(b) might do so for a suitable $\Gamma(s)$, thus resulting in a stable feedback interconnection ${H}^{\prime}_{1}\#{H}^{\prime}_{2}$.
+Notice that this dynamic loop-shifting approach based on the dynamic transfer function $\Gamma(s)$ effectively corresponds to the concept of frequency-dependent (dynamic) passivity indices[18, 19], which, are generally less conservative than static passivity indices.
 Among the four closed-loop transfer matrices in 2 relevant to internal feedback stability, it is immediately evident that only the upper-left transfer matrix mapping from $w_{1}$ to $y_{1}$ remains equivalent between the original system in Fig.
@@
 By swapping the feedforward and feedback of $\Gamma(s)$ in Fig. 1(b) and comparing the closed-loop transfer function from $w_{2}$ to $y_{2}$, similar stability conclusions can be made.
+Remark 2.
+Previous dynamic loop-shifting techniques based on the transfer matrix $\Gamma(s)$ are purely mathematical tools for establishing passivity properties and analytically proving closed-loop stability.
+They do not involve any hardware changes or modifications to the physical control system.
 3 Power System Model
@@
 We consider a Kron-reduced [21], balanced network with $n\in\mathbb{N}$ converter nodes, denoted by $\{1,...,n\}$, where the dynamic small-signal model (in the global per unit system) of the line between node $i$ and node $j$ ($i,j\in\{1,...,n\}$) is given in the frequency domain as
-$\begin{bmatrix}\Delta i_{\mathrm{d},ij}(s)\\ \Delta i_{\mathrm{q},ij}(s)\end{bmatrix}=y_{ij}(s)\left(\begin{bmatrix}\Delta v_{\mathrm{d},i}(s)\\ \Delta v_{\mathrm{q},i}(s)\end{bmatrix}-\begin{bmatrix}\Delta v_{\mathrm{d},j}(s)\\ \Delta v_{\mathrm{q},j}(s)\end{bmatrix}\right),$ (4)
+[…]
 where $\Delta i_{\mathrm{dq},ij}=\left[\Delta i_{\mathrm{d},ij}\,\,\Delta i_{\mathrm{q},ij}\right]^{\top}$ is the current vector from node $i$ to node $j$ in the global $\mathrm{dq}$ coordinates with a constant nominal rotating frequency $\omega_{0}$, $\Delta v_{\mathrm{dq},i}=\left[\Delta v_{\mathrm{d},i}\,\,\Delta v_{\mathrm{q},i}\right]^{\top}$ is the voltage vector of node $i$, and $y_{ij}(s)$ is a $2\times 2$ transfer matrix encoding the small-signal dynamics of the line, i.e.,
@@
 where $b_{ij}=\tfrac{1}{l_{ij}}$ is the line susceptance, and $\rho_{ij}=\tfrac{r_{ij}}{l_{ij}}$ is the resistance-inductance ratio of the line $ij$ which is assumed to be small and uniform (i.e., $\rho_{ij}=\rho\ll 1,\,\forall\,i,j\in\{1,...,n\}$) throughout the dominantly inductive transmission network.
-Notice that $b_{ij}=0$ and $\rho_{ij}=0$ if there is no line.
+If there is no line, we consider $b_{ij}=0$ and $\rho_{ij}=0$.
+As is standard in small-signal analysis, the network model assumes fixed RL-line parameters and a constant $\omega_{0}$ in the applied $\mathrm{dq}$ transformation, and is therefore valid only in the small-signal regime around the nominal operating frequency.
 […]
@@
 The time-domain bus voltages are modeled in polar coordinates as $v_{\mathrm{d},i}(t)\coloneqq|v|_{i}(t)\cos{\delta_{i}(t)}$ and $v_{\mathrm{q},i}(t)\coloneqq|v|_{i}(t)\sin{\delta_{i}(t)}$ with magnitude $|v|_{i}(t)$ and relative angle $\delta_{i}(t)=\theta_{i}(t)-\theta_{0}(t)$ in SI units, where $\theta_{i}(t)$ is the voltage angle at bus $i\in\{1,...,n\}$, and $\theta_{0}(t)$ the Park transformation angle of the global $\mathrm{dq}$ coordinates, which can be expressed as
-$\begin{split}|v|_{i}(t)=\sqrt{v_{\mathrm{d},i}(t)^{2}+v_{\mathrm{q},i}(t)^{2}},\quad\delta_{i}(t)=\arctan\left(\tfrac{v_{\mathrm{q},i}(t)}{v_{\mathrm{d},i}(t)}\right).\end{split}$ (8)
+$\begin{split}\hskip-5.69054pt|v|_{i}(t)\hskip-1.99168pt=\hskip-1.99168pt\sqrt{\hskip-1.99168ptv_{\mathrm{d},i}(t)^{2}\hskip-1.99168pt+\hskip-1.99168ptv_{\mathrm{q},i}(t)^{2}},\quad\delta_{i}(t)\hskip-1.99168pt=\arctan\hskip-1.99168pt\left(\tfrac{v_{\mathrm{q},i}(t)}{v_{\mathrm{d},i}(t)}\right)\hskip-1.42262pt.\hskip-5.69054pt\end{split}$ (9)
 We consider the active and reactive branch powers
@@
 $\begin{split}p_{i}(t)&=v_{\mathrm{d},i}(t)i_{\mathrm{d},i}(t)+v_{\mathrm{q},i}(t)i_{\mathrm{q},i}(t)\\ q_{i}(t)&=-v_{\mathrm{d},i}(t)i_{\mathrm{q},i}(t)+v_{\mathrm{q},i}(t)i_{\mathrm{d},i}(t).\end{split}$ (11)
-By linearizing 8 and 10, transforming them into the frequency domain, and performing some analytical computations (see Appendix 7 for details), we obtain
-$\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta p(s)\\ \Delta q(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta p_{1}(s)\\ \Delta q_{1}(s)\\ \vdots\\ \Delta p_{n}(s)\\ \Delta q_{n}(s)\end{bmatrix}}}=\underset{\eqqcolon\,{N}(s)}{\underbrace{\begin{bmatrix}N_{11}(s)&\cdots&N_{1n}(s)\\ \vdots&\ddots&\vdots\\ N_{n1}(s)&\cdots&N_{nn}(s)\end{bmatrix}}}\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta\delta(s)\\ \Delta|v|_{\mathrm{n}}(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta\delta_{1}(s)\\ \Delta{|v|}_{\mathrm{n},1}(s)\\ \vdots\\ \Delta\delta_{n}(s)\\ \Delta{|v|}_{\mathrm{n},n}(s)\end{bmatrix}}}$ (11)
+By linearizing 9 and 11, transforming them into the frequency domain, and performing some analytical computations (see Appendix 7 for details), we obtain
+$\hskip-2.84526pt\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta p(s)\\ \Delta q(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta p_{1}(s)\\ \Delta q_{1}(s)\\ \vdots\\ \Delta p_{n}(s)\\ \Delta q_{n}(s)\end{bmatrix}}}\hskip-1.42262pt=\hskip-1.42262pt\underset{\eqqcolon\,{N}(s)}{\underbrace{\begin{bmatrix}N_{11}(s)&\cdots&N_{1n}(s)\\ \vdots&\ddots&\vdots\\ N_{n1}(s)&\cdots&N_{nn}(s)\end{bmatrix}}}\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta\delta(s)\\ \Delta|v|_{\mathrm{n}}(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta\delta_{1}(s)\\ \Delta{|v|}_{\mathrm{n},1}(s)\\ \vdots\\ \Delta\delta_{n}(s)\\ \Delta{|v|}_{\mathrm{n},n}(s)\end{bmatrix}}}\hskip-2.84526pt$ (12)
 with the $2\times 2$ transfer matrix blocks $N_{ii}(s)$ and $N_{ij}(s)$
-$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}1&\rho+\tfrac{s}{\omega_{0}}\\ -(\rho+\tfrac{s}{\omega_{0}})&1\end{bmatrix}\\ &\quad\quad\quad+\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}-1&\rho\\ \rho&1\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}-1&-(\rho+\tfrac{s}{\omega_{0}})\\ (\rho+\tfrac{s}{\omega_{0}})&-1\end{bmatrix},\end{split}$ (12)
+[…]
 where $\Delta{|v|}_{\mathrm{n},i}(s)\coloneqq\tfrac{\Delta|v|_{i}(s)}{|v|_{0,i}}$ is the voltage magnitude at bus $i$ normalized at the steady-state $|v|_{0,i}$.
-Moreover, to derive 12, as in any power system small-signal model, we have assumed a small steady-state angle difference $\delta_{0,i}\approx\delta_{0,j}$, which is standard and justified since thermal limitations for transmission lines preclude large angle differences [20].
+Moreover, to derive 13, as in any power system small-signal model, we have assumed a small steady-state angle difference $\delta_{0,i}\approx\delta_{0,j}$, which is standard and justified since thermal limitations for transmission lines preclude large angle differences [22].
 For a dominantly inductive transmission network with $\rho\ll 1$, the standard approximation of the line dynamics matrix
 $\begin{split}M(s)&=\tfrac{1}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}1&\rho+\tfrac{s}{\omega_{0}}\\ -(\rho+\tfrac{s}{\omega_{0}})&1\end{bmatrix}\end{split}$ (14)
-appearing in $N_{ii}(s)$ and $N_{ij}(s)$ in 12 is
-$\begin{split}M_{1}(s)&=\tfrac{1}{1+\tfrac{s^{2}}{\omega_{0}^{2}}}\begin{bmatrix}1&\tfrac{s}{\omega_{0}}\\ -\tfrac{s}{\omega_{0}}&1\end{bmatrix},\end{split}$ (14)
+appearing in $N_{ii}(s)$ and $N_{ij}(s)$ in 13 is
+[…]
 i.e., losses are entirely neglected ($\rho=0$). Here we put forward a novel and more accurate approximation for $M(s)$, namely
@@
 Indeed, a straightforward calculation comparing the Euclidean norms of the matrix distances for $s=\mathrm{j}\omega$ reveals that $\forall\omega\geq 0$
-$\tfrac{||M(\mathrm{j}\omega)-M_{2}(\mathrm{j}\omega)||_{F}}{||M(\mathrm{j}\omega)-M_{1}(\mathrm{j}\omega)||_{F}}=\sqrt{\tfrac{1+\tfrac{\omega^{4}}{\omega_{0}^{4}}-2\tfrac{\omega^{2}}{\omega_{0}^{2}}}{1+\tfrac{\omega^{4}}{\omega_{0}^{4}}+6\tfrac{\omega^{2}}{\omega_{0}^{2}}+\rho^{2}\left(1+\tfrac{\omega^{2}}{\omega_{0}^{2}}\right)}}<1,$ (16)
+[…]
 that is, $M_{2}$ is a strictly better approximation for $\rho>0$.
 For $\rho\rightarrow 0$, it can be shown that both approximations are consistent in terms of recovering $M$ asymptotically.
-Using the approximation in 15, we can therefore eliminate the less dominant terms of the off-diagonal matrix elements in 12 as stated below.
+Using the approximation 16, we can eliminate the less dominant terms of the off-diagonal matrix elements in 13 as stated below, which allows to exploit symmetry properties of the network dynamics in the stability proof in Section 4.3.
+However, since 16 is valid only for $\rho\ll 1$, extending our results to more general R/X ratios remains a topic for future work.
 Definition 4 (Dynamic Small-Signal Network Model).
-The transfer matrix blocks for the dynamic (i.e., $s\neq 0$, $|v|_{0,i}\neq|v|_{0,j}$) small-signal network model in 11 used to derive the stability certificates in Section 4 are given by:
-$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}1&\tfrac{s}{\omega_{0}}\\ -\tfrac{s}{\omega_{0}}&1\end{bmatrix}\\ &\quad\quad\quad+\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}\begin{bmatrix}-1&-\tfrac{s}{\omega_{0}}\\ \tfrac{s}{\omega_{0}}&-1\end{bmatrix}.\end{split}$ (17)
-By setting $s=0$ in 17, we derive the quasi-stationary network model, which is employed in a similar form in [9]:
+The transfer matrix blocks for the dynamic (i.e., $s\neq 0$, $|v|_{0,i}\neq|v|_{0,j}$) small-signal network model in 12 used to derive the stability certificates in Section 4 are given by:
+[…]
+By setting $s=0$ in 18, we derive the quasi-stationary network model, which is employed in a similar form in [10]:
 Definition 5 (Network-Simplification Level 1).
-The transfer matrix blocks for the quasi-stationary (i.e., $s=0$, $|v|_{0,i}\neq|v|_{0,j}$) small-signal network model in 11 are given by:
-$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{1}{1+\rho^{2}}\begin{bmatrix}|v|_{0,i}|v|_{0,j}&0\\ 0&2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}-1&0\\ 0&-1\end{bmatrix}.\end{split}$ (18)
-To further simplify 18, we set $|v|_{0,i}=|v|_{0,j}=|v|_{0}$, yielding the zero-power flow network model similar to [10]:
+The transfer matrix blocks for the quasi-stationary (i.e., $s=0$, $|v|_{0,i}\neq|v|_{0,j}$) small-signal network model in 12 are given by:
+[…]
+To further simplify 19, we set $|v|_{0,i}=|v|_{0,j}=|v|_{0}$, yielding the zero-power flow network model similar to [11]:
 Definition 6 (Network-Simplification Level 2).
-The transfer matrix blocks for the zero-power flow (i.e., $s=0$, $|v|_{0,i}=|v|_{0,j}=|v|_{0}$) small-signal network model in 11 are given by:
-$\begin{split}N_{ii}(s)&=\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0}^{2}}{1+\rho^{2}}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\\ N_{ij}(s)&=b_{ij}\tfrac{|v|_{0}^{2}}{1+\rho^{2}}\begin{bmatrix}-1&0\\ 0&-1\end{bmatrix}.\end{split}$ (19)
+The transfer matrix blocks for the zero-power flow (i.e., $s=0$, $|v|_{0,i}\hskip-1.42262pt=\hskip-1.42262pt|v|_{0,j}\hskip-1.42262pt=\hskip-1.42262pt|v|_{0}$) small-signal network model in 12 are given by:
+[…]
+Remark 3.
+For $\rho=0$, the network model in [10] corresponds to 19 in Definition 5, while the model in [11] aligns with 20 in Definition 6.
+Moreover, other works [9, 8] focus solely on SISO frequency dynamics.
+Hence, our dynamic network model in 18 offers the most detailed representation in literature which is compatible with theoretical stability certificates.
 […]
 Figure 3: Basic implementation of a three-phase GFM VSC.
-Remark 2.
-For $\rho=0$, the network model in [9] corresponds to 18 in Definition 5, while the model in [10] aligns with 19 in Definition 6.
-Moreover, other works [8, 7] focus solely on SISO frequency dynamics.
-Hence, our dynamic network model in 17 offers the most detailed representation in literature which is compatible with theoretical stability certificates.
 3.2 Grid-Forming Converter Dynamics
 Fig.
-3 shows the basic implementation of a GFM three-phase VSC connected to the power network in Fig.
+3 shows the implementation of a GFM three-phase VSC [^1] A simple three-phase VSC implementation is used here for illustration.
+However, our results are not limited to this specific converter architecture and can be extended to other filter configurations and inner-loop structures, provided the time-scale separation discussed in Remark 4 holds.
+connected to the power network in Fig.
 2.
-The linearized small-signal dynamics (in the global per unit system) of the $i$-th VSC ($i\in\{1,...,n\}$) can be represented by the $2\times 2$ transfer matrix $D_{i}(s)$, mapping from the active and reactive power injections $\Delta p_{i}(s)$ and $\Delta q_{i}(s)$ to the imposed frequency and voltage magnitude $\Delta\omega_{i}(s)=\Delta\delta_{i}(s)s$ and $\Delta|v|_{i}(s)$, i.e.,
+The linearized small-signal dynamics (in the global per unit system) of the $i$-th VSC ($i\in\{1,...,n\}$) can be represented by the $2\times 2$ transfer matrix $D_{i}(s)$, mapping from active and reactive power injections $\Delta p_{i}(s)$ and $\Delta q_{i}(s)$ to the imposed frequency and voltage magnitude $\Delta\omega_{i}(s)=\Delta\delta_{i}(s)s$ and $\Delta|v|_{i}(s)$, i.e.,
 $-\begin{bmatrix}\Delta\omega_{i}(s)\\ \Delta|v|_{i}(s)\end{bmatrix}=D_{i}(s)\begin{bmatrix}\Delta p_{i}(s)\\ \Delta q_{i}(s)\end{bmatrix}$. (21)
 The most prevalent GFM control strategy is filtered droop control [23].
-Assuming time-scale separation, we can neglect the internal dynamics of the VSC (see Appendix 8 for a detailed derivation) and consider $D_{i}(s)$ as in the following definition.
+Assuming time-scale separation as elaborated in Remark 4, we can neglect the internal dynamics of the VSC (see Appendix 8 for a derivation) and consider $D_{i}(s)$ as in the following definition.
 Of course, all our developments will be evaluated on the full VSC model in Section 5.
 Definition 7 (Dynamic Small-Signal Converter Model).
-The small-signal converter dynamics 20 used to derive the stability certificates in Section 4 are given by the droop controller
+The small-signal converter dynamics 21 used to derive the stability certificates in Section 4 are given by the droop controller
 $D_{i}(s)=\begin{bmatrix}\tfrac{d_{\mathrm{p},i}}{\tau_{\mathrm{p},i}s+1}&0\\ 0&\tfrac{d_{\mathrm{q},i}}{\tau_{\mathrm{q},i}s+1}\end{bmatrix}$, (22)
 where $d_{\mathrm{p},i}\in\mathbb{R}_{>0}$ and $d_{\mathrm{q},i}\in\mathbb{R}_{>0}$ are the active and reactive power droop gains, and $\tau_{\mathrm{p},i}\in\mathbb{R}_{\geq 0}$ and $\tau_{\mathrm{q},i}\in\mathbb{R}_{\geq 0}$ are the active and reactive power low-pass filter time constants.
-Finally, if we extend 20 to include the dynamics of all $n$ converters, we obtain the following $2n\times 2n$ transfer matrix
-$-\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta\omega(s)\\ \Delta|v|(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta\omega_{1}(s)\\ \Delta{|v|}_{1}(s)\\ \vdots\\ \Delta\omega_{n}(s)\\ \Delta{|v|}_{n}(s)\end{bmatrix}}}=\underset{\eqqcolon\,D(s)}{\underbrace{\begin{bmatrix}D_{1}(s)&0_{2\times 2}&\dots&0_{2\times 2}\\ 0_{2\times 2}&D_{2}(s)&\dots&0_{2\times 2}\\ \vdots&\vdots&\ddots&\vdots\\ 0_{2\times 2}&0_{2\times 2}&\dots&D_{n}(s)\end{bmatrix}}}\underset{\eqqcolon\,{\tiny\begin{bmatrix}\Delta p(s)\\ \Delta q(s)\end{bmatrix}}}{\underbrace{\begin{bmatrix}\Delta p_{1}(s)\\ \Delta q_{1}(s)\\ \vdots\\ \Delta p_{n}(s)\\ \Delta q_{n}(s)\end{bmatrix}}}.$ (22)
-$\mathcal{D}_{0}(s)$$\mathcal{N}_{0}(s)$--$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|v|\end{bmatrix}$$\begin{bmatrix}\Delta p\\ \Delta q\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|v|_{\mathrm{d}}\end{bmatrix}$device dynamicsnetwork dynamics--$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$
+Finally, if we extend 21 to include the dynamics of all $n$ converters, we obtain the following $2n\times 2n$ transfer matrix
+[…]
+Remark 4.
+From a power electronics perspective, three distinct dynamic time scales are typically considered (see Fig. 5):
+  • (i)
+    fast inner VSC control loops,
+  • (ii)
+    intermediate network dynamics, and
+  • (iii)
+    slow outer VSC control loops.
+These layers must be sufficiently separated in this order; otherwise, standard assumptions for control design and model reduction break down.
+Consistent with [9, 12], in this work, we assume that inner-loop and network dynamics are well separated, while a possible overlap between network and outer-loop dynamics is treated as a separate problem.
+$\mathcal{D}_{0}(s)$$\mathcal{N}_{0}(s)$-$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|v|\end{bmatrix}$$\begin{bmatrix}\Delta p\\ \Delta q\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|v|_{\mathrm{d}}\end{bmatrix}$device dynamicsnetwork dynamics-$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$
 Figure 4: Closed-loop power system dynamics where $\Delta p_{\mathrm{d}}$ and $\Delta q_{\mathrm{d}}$ are the active and reactive power disturbances, $\Delta\omega_{\mathrm{d}}$ is the frequency disturbance, and $\Delta|{v}|_{\mathrm{d}}$ is the voltage magnitude disturbance.
+1msphysicsconrolinner convertercontrol (current+ voltage loop)outer convertercontrol (active &reactive power loopsnetworkdynamicstimescales of interest10ms100ms
+Figure 5: Time scales of interest for our stability study, adapted from [12].
 3.3 Dynamic Power System Model
-The closed-loop power system dynamics are modeled as the feedback interconnection of the converter device dynamics in 22 and the network dynamics in 11 as illustrated in Fig.
-4.
+The closed-loop power system dynamics are modeled as the feedback interconnection of the converter device dynamics in 23 and the network dynamics in 12 as illustrated in Fig.
+4222An extension to heterogeneous device allocations with synchronous generators, grid-forming and -following converters will be part of future work..
 We focus on the stability of the bus frequency and voltage magnitudes, given by $\Delta\omega_{i}(s)=\Delta\delta_{i}(s)s$ and $\Delta|v|_{i}(s)$ for $i\in\{1,...,n\}$ and consider these quantities as interconnection signals between the device and network dynamics in Fig.
@@
 In the next section, we establish internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$ under certain decentralized parametric conditions.
-$c_{i,\rho}$ Definition $c_{i,\rho}$ Definition
-$c_{1,\rho}$ $({1+\rho^{2}})/({5+2\rho^{2}})$ $c_{8,\rho}$ $2\rho\omega_{0}(15|v|_{\mathrm{max}}^{2}+4)(10\rho^{2}|v|_{\mathrm{max}}^{2}+25|v|_{\mathrm{max}}^{2}+4)$
-$c_{2,\rho}$ ${(1+\rho^{2})^{2}}/({6\rho})$ $c_{9,\rho}$ $10\omega_{0}\rho^{4}|v|_{\mathrm{max}}^{2}+95\omega_{0}\rho^{2}|v|_{\mathrm{max}}^{2}+16\omega_{0}\rho^{2}+25\omega_{0}|v|_{\mathrm{max}}^{2}$
-$c_{3,\rho}$ $2\rho(1+\rho^{2})$ $c_{10,\rho}$ $5\omega_{0}(1+\rho^{2})(20\rho^{3}|v|_{\mathrm{max}}^{2}+80\rho|v|_{\mathrm{max}}^{2}+16\rho)$
-$c_{4,\rho}$ $((2\rho^{2}-5)+\sqrt{(5-2\rho^{2})^{2}+16\rho^{2}})/4\rho$ $c_{11,\rho}$ $50\rho\omega_{0}(1+\rho^{2})^{2}$
-$c_{5,\rho}$ ${(1+\rho^{2})}/({10\rho^{2}|v|_{\mathrm{max}}^{2}+25|v|_{\mathrm{max}}^{2}+4)}$ $c_{12,\rho}$ ${(\rho^{2}+1)^{2}}/({6\rho|v|_{\mathrm{max}}^{2}})$
-$c_{6,\rho}$ ${5(1+\rho^{2})^{2}}/({30\rho|v|_{\mathrm{max}}^{2}+8\rho)}$ $c_{13,\rho}$ $-2\rho|v|_{\mathrm{max}}^{4}$
-$c_{7,\rho}$ ${5(1+\rho^{2})}/{4}$ $c_{14,\rho}$ $|v|_{\mathrm{max}}^{2}(2\rho^{2}-5)$
-Table 1: Definition of the quantities $c_{l,\rho}$ for $l=1,...,14$ in the conditions 24 and 25 depending on the parameter $\rho$.
-(a) 2D plot with $\rho=0.3$.
-(b) 3D plot with variable $\rho$.
-Figure 5: Graphical illustration of the stability conditions in 24, where $\alpha_{\mathrm{p},i}=\tfrac{d_{\mathrm{p},i}}{\omega_{0}}|v|_{\mathrm{max}}^{2}\sum_{j\neq i}^{n}b_{ij}$ and $\tilde{\tau}_{\mathrm{p},i}=\tau_{\mathrm{p},i}\omega_{0}$,.
-(a) 2D plot with $\rho=0.3$.
-(b) 3D plot with variable $\rho$.
-Figure 6: Graphical illustration of the stability conditions in 25, where $\alpha_{\mathrm{q},i}=\tfrac{d_{\mathrm{q},i}}{|v|_{0,i}}\sum_{j\neq i}^{n}b_{ij}$ and $\tilde{\tau}_{\mathrm{q},i}=\tau_{\mathrm{q},i}\omega_{0}$.
+$\hskip-2.84526ptc_{i,\rho}\hskip-2.84526pt$ Definition $\hskip-2.84526ptc_{i,\rho}\hskip-2.84526pt$ Definition
+$\hskip-2.84526ptc_{1,\rho}\hskip-2.84526pt$ $({1+\rho^{2}})/({5+2\rho^{2}})$ $c_{6,\rho}\hskip-2.84526pt$ ${(1+\rho^{2})}/({2\rho^{2}|v|_{\mathrm{max}}^{2}+5|v|_{\mathrm{max}}^{2})}$
+$\hskip-2.84526ptc_{2,\rho}\hskip-2.84526pt$ ${(1+\rho^{2})^{2}}/({6\rho})$ $c_{7,\rho}$ ${(\rho^{2}+1)^{2}}/({6\rho|v|_{\mathrm{max}}^{2}})$
+$\hskip-2.84526ptc_{3,\rho}\hskip-2.84526pt$ $2\rho(1+\rho^{2})$ $c_{8,\rho}$ $2\rho|v|_{\mathrm{max}}^{4}$
+$\hskip-2.84526ptc_{4,\rho}\hskip-2.84526pt$ $((2\rho^{2}-5)+\sqrt{(5-2\rho^{2})^{2}+16\rho^{2}})/4\rho$ $c_{9,\rho}$ $|v|_{\mathrm{max}}^{2}(2\rho^{2}-5)$
+$\hskip-2.84526ptc_{5,\rho}\hskip-2.84526pt$ ${5(1+\rho^{2})}/{4}$
+Table 1: Definition of the quantities $c_{l,\rho}$ for l=1,,9l=1,...,9 in the conditions 25 and 26 depending on the parameter $\rho$.
 4 Decentralized Stability Conditions
-By applying dynamic loop-shifting techniques and passivity theory (see Section 4.2 for details), we derive parametric decentralized stability conditions that serve as local tuning rules for the dynamic droop control 21 of each VSC $i\in\{1,\dots,n\}$, ensuring the internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$ in Fig.
+In this section, we first introduce the decentralized stability conditions, which provide practical guidelines for local converter tuning and decentralized stability assessment.
+We then present our main result, which guarantees internal feedback stability of the closed-loop system, assuming the previously stated stability conditions are satisfied.
+Notably, these conditions naturally emerge from the proof of the main result.
+4.1 Decentralized Stability Conditions
+By applying dynamic loop-shifting techniques and passivity theory (see Section 4.3 for details), we derive parametric decentralized stability conditions that serve as local tuning rules for the dynamic droop control 22 of each VSC $i\in\{1,\dots,n\}$, ensuring the internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$ in Fig.
 4.
@@
 Crucially, the conditions are local and entirely independent of the control parameters of other VSCs, making them directly applicable for scalable and decentralized stability assessment, as well as for local controller tuning and grid-code design.
+Beyond that, the decentralized stability certificates support heterogeneous device-level controllers by allowing each converter to be designed with independent droop gains and time constants, without requiring uniform parameters across the system, as is the case in, e.g., [9].
 The decentralized stability conditions for each VSC $i$ can be categorized into decoupled constraints on the active power-frequency droop control and the reactive power-voltage droop control.
-For each control scheme, these conditions can be visualized in either a two-dimensional (for fixed $\rho$) or a three-dimensional (for variable $\rho$) coordinate system, as shown in Figures 5 and 6.
+For each control scheme, these conditions can be visualized in either a two-dimensional (for fixed $\rho$) or a three-dimensional (for variable $\rho$) coordinate system, as shown in Figures 6 and 7.
 The coordinate axes represent scaled versions of the local droop gains and time constants (in 2D) and the global network parameter $\rho$ (in 3D), with typical value ranges.
@@
 As illustrated in Fig.
-5, for small values of $\rho$, closed-loop stability is ensured when the droop gain $d_{\mathrm{p},i}$ and/or the self-susceptance $\sum_{j\neq i}^{n}b_{ij}$ are sufficiently small, irrespective of the time constant $\tau_{\mathrm{p},i}$.
+6, for small values of $\rho$, closed-loop stability is ensured when the droop gain $d_{\mathrm{p},i}$ and/or the self-susceptance $\sum_{j\neq i}^{n}b_{ij}$ are sufficiently small, irrespective of the time constant $\tau_{\mathrm{p},i}$.
 Notably, this suggests potential instability issues in future transmission systems with a high density of devices interconnected by short lines, which correspond to large values of $\sum_{j\neq i}^{n}b_{ij}$.
@@
 In particular, in contrast to $\tau_{\mathrm{p},i}$, a larger $\tau_{\mathrm{q},i}$ can be stabilizing.
-Finally, from the 3D plots in Figures 5 and 6, it becomes apparent how an increasing $\rho$ allows for larger local droop gains of both the active and reactive power droop control.
-In what follows, we provide an algebraic parametrization of the conditions in Figures 5 and 6 (a detailed derivation is provided in Section 4.2).
+Finally, from the 3D plots in Figures 6 and 7, it becomes apparent how an increasing $\rho$ allows for larger local droop gains of both the active and reactive power droop control.
+(a) 2D plot with $\rho=0.3$.
+(b) 3D plot with variable $\rho$.
+Figure 6: Graphical illustration of the stability conditions in 25, where $\alpha_{\mathrm{p},i}\hskip-1.42262pt=\hskip-1.42262pt\tfrac{d_{\mathrm{p},i}}{\omega_{0}}|v|_{\mathrm{max}}^{2}\sum_{j\neq i}^{n}b_{ij}$ and $\tilde{\tau}_{\mathrm{p},i}\hskip-1.42262pt=\hskip-1.42262pt\tau_{\mathrm{p},i}\omega_{0}$ (in rad).
+(a) 2D plot with $\rho=0.3$.
+(b) 3D plot with variable $\rho$.
+Figure 7: Graphical illustration of the stability conditions in 26, where $\alpha_{\mathrm{q},i}=\tfrac{d_{\mathrm{q},i}}{|v|_{0,i}}\sum_{j\neq i}^{n}b_{ij}$ and $\tilde{\tau}_{\mathrm{q},i}=\tau_{\mathrm{q},i}\omega_{0}$ (in rad).
+In what follows, we provide an algebraic parametrization of the conditions in Figures 6 and 7 (a detailed derivation is provided in Section 4.3).
 Namely, for the active power-frequency droop control, we require
@@
 $\begin{split}\alpha_{\mathrm{p},i}&<c_{4\rho},\end{split}$ (25d)
-where $\alpha_{\mathrm{p},i}=\tfrac{d_{\mathrm{p},i}}{\omega_{0}}|v|_{\mathrm{max}}^{2}\sum_{j\neq i}^{n}b_{ij}$, $\tilde{\tau}_{\mathrm{p},i}=\tau_{\mathrm{p},i}\omega_{0}$, and $c_{l\rho}$ for $l=\{1,2,3,4\}$ are quantities depending on $\rho$ as defined in Table 1.
+where $\alpha_{\mathrm{p},i}\hskip-1.42262pt=\hskip-1.42262pt\tfrac{d_{\mathrm{p},i}}{\omega_{0}}|v|_{\mathrm{max}}^{2}\sum_{j\neq i}^{n}b_{ij}$, $\tilde{\tau}_{\mathrm{p},i}\hskip-1.42262pt=\hskip-1.42262pt\tau_{\mathrm{p},i}\omega_{0}$, and $c_{l\rho}$ for $l\hskip-1.42262pt=\hskip-1.42262pt\{1,2,3,4\}$ are quantities depending on $\rho$ as defined in Table 1.
 For the reactive power-voltage droop control, we require
-$\begin{split}\alpha_{\mathrm{q},i}&<c_{5,\rho}(5+10\rho\tilde{\tau}_{\mathrm{q},i})\end{split}$ (25a)
-$\begin{split}\alpha_{\mathrm{q},i}&<c_{6,\rho}(2\rho+\tilde{\tau}_{\mathrm{q},i})\end{split}$ (25b)
-$\begin{split}\alpha_{\mathrm{q},i}&<c_{7,\rho}\end{split}$ (25c)
-$\begin{split}0&<\alpha_{\mathrm{q},i}^{2}c_{8\rho}-\alpha_{\mathrm{q},i}(5\tilde{\tau}_{\mathrm{q},i}(1+\rho^{2})c_{9,\rho}+c_{10,\rho})\\ &\quad\quad\quad+c_{11,\rho}(\tilde{\tau}_{\mathrm{q},i}^{2}(1+\rho^{2})+2\rho\tilde{\tau}_{\mathrm{q},i}+1)\end{split}$ (25d)
-$\begin{split}\alpha_{\mathrm{q},i}&<c_{12,\rho}\tilde{\tau}_{\mathrm{q},i}\end{split}$ (25e)
-$\begin{split}0&<\alpha_{\mathrm{q},i}^{2}c_{13,\rho}+\alpha_{\mathrm{q},i}\tilde{\tau}_{\mathrm{q},i}c_{14,\rho}+2\tilde{\tau}_{\mathrm{q},i}^{2}\rho\end{split}$ (25f)
+$\begin{split}\alpha_{\mathrm{q},i}&<c_{5,\rho}\end{split}$ (26a)
+$\begin{split}\alpha_{\mathrm{q},i}&<c_{6,\rho}2\rho\tilde{\tau}_{\mathrm{q},i}\end{split}$ (26b)
+$\begin{split}\alpha_{\mathrm{q},i}&<c_{7,\rho}\tilde{\tau}_{\mathrm{q},i}\end{split}$ (26c)
+$\begin{split}0&<\alpha_{\mathrm{q},i}^{2}c_{8,\rho}+\alpha_{\mathrm{q},i}\tilde{\tau}_{\mathrm{q},i}c_{9,\rho}+2\tilde{\tau}_{\mathrm{q},i}^{2}\rho\end{split}$ (26d)
 where $\alpha_{\mathrm{q},i}=\tfrac{d_{\mathrm{q},i}}{|v|_{0,i}}\sum_{j\neq i}^{n}b_{ij}$, $\tilde{\tau}_{\mathrm{q},i}=\tau_{\mathrm{q},i}\omega_{0}$, and $c_{l\rho}$ for $l=\{5,...,14\}$ are quantities depending on $\rho$ as defined in Table 1.
 Remark 5 (Plug-and-Play Operation).
-The conditions in 24 and 25 are designed for the scenario where all $n$ converter nodes are populated, ensuring system-level stability during disconnection and re-connection of VSCs.
-Specifically, removing any VSC decreases the self susceptance $\sum_{j\neq i}^{n}b_{ij}$ for each connected VSC $i$ after Kron reduction and reindexing, thereby guaranteeing that 24 and 25 remain locally satisfied.
-[…]
-Figure 7: Structural overview of the proof of Theorem 2.
+The conditions in 25 and 26 are designed for the scenario where all $n$ converter nodes are populated, ensuring system-level stability during disconnection and re-connection of VSCs.
+Specifically, removing any VSC decreases the self susceptance $\sum_{j\neq i}^{n}b_{ij}$ for each connected VSC $i$ after Kron reduction and reindexing, thereby guaranteeing that 25 and 26 remain locally satisfied.
 4.2 Main Result
-The conditions 24 and 25 are derived in Section 4.2, which proves our main result stated in the theorem below:
+The conditions 25 and 26 are derived in Section 4.3, which proves our main result stated in the theorem below:
 Theorem 2 (Internal Feedback Stability of the Closed-Loop System).
-Consider the device and network dynamics $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$ in 23, where the network model is given by 17.
-Let the conditions in 24 and 25 hold for each VSC $i\in\{1,...,n\}$.
+Consider the device and network dynamics $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$ in 24, where the network model is given by 18.
+Let the conditions in 25 and 26 hold for each VSC $i\in\{1,...,n\}$.
 Then, the closed-loop system $\mathcal{D}_{0}\#\mathcal{N}_{0}$ in Fig.
@@
 Corollary 1 (Simplified Network Models).
-For the quasi-stationary network dynamics $\mathcal{N}_{0}(s)$ with the network model in 18, where $s=0$ and $|v|_{0,i}\neq|v|_{0,j}$, the stability result of Theorem 2 holds under the relaxed algebraic conditions [^1] These conditions structurally align with the conditions derived in [9].
-$\begin{split}\alpha_{\mathrm{q},i}<\tfrac{5(1+\rho^{2})}{4}=c_{7,\rho}\quad\text{and}\quad 0<\tau_{\mathrm{p},i}\end{split}$ (26)
+For the quasi-stationary network dynamics $\mathcal{N}_{0}(s)$ with the network model in 19, where $s=0$ and $|v|_{0,i}\neq|v|_{0,j}$, the stability result of Theorem 2 holds under the relaxed algebraic conditions
+$\begin{split}\alpha_{\mathrm{q},i}<\tfrac{5(1+\rho^{2})}{4}=c_{5,\rho}\quad\text{and}\quad 0<\tau_{\mathrm{p},i}\end{split}$ (27)
 for each VSC $i\in\{1,...,n\}$.
-Further, for the zero-power-flow network dynamics $\mathcal{N}_{0}(s)$ with the network model in 19, where $s=0$ and $|v|_{0,i}=|v|_{0,j}$, internal feedback stability is guaranteed without imposing any additional VSC conditions.
-The proof of Corollary 1 is provided in Appendix 9 and follows a similar reasoning as the proof of Theorem 2 in Section 4.2, but applies the simplified network models in 18 and 19, respectively.
-It becomes apparent how the conditions in 24 and 25 are a subset of the conditions in 26.
+Further, for the zero-power-flow network dynamics $\mathcal{N}_{0}(s)$ with the network model in 20, where $s=0$ and $|v|_{0,i}=|v|_{0,j}$, internal feedback stability is guaranteed without imposing any additional VSC conditions.
+The proof of Corollary 1 is provided in Appendix 9 and follows a similar reasoning as the proof of Theorem 2 in Section 4.3.
+It becomes apparent how the conditions in 25 and 26 are a subset of the conditions in 27.
+The latter structurally align with the conditions derived in [10].
 4.3 Proof of Theorem 2
@@
 We then analyze the stability of the closed-loop system as in Fig.
-8(a), where
+9(a), where
 $\begin{split}\mathcal{D}(s)&\coloneqq D(s)\cdot\mathrm{diag}(1,\tfrac{s}{|v|_{0,1}},\dots,1,\tfrac{s}{|v|_{0,n}})\\ \mathcal{N}(s)&\coloneqq N(s)\cdot\mathrm{diag}(\tfrac{1}{s},\tfrac{1}{s},\dots,\tfrac{1}{s},\tfrac{1}{s}).\end{split}$ (28)
@@
 Given that $\mathcal{D}(s)$ and $\mathcal{N}(s)$ in Fig.
-8(a) do not satisfy the passivity conditions in Theorem 1 (i.e., it can be shown that $\mathcal{N}(s)$ is not passive), we resort to dynamic loop-shifting techniques as presented in Section 2.3.
+9(a) do not satisfy the passivity conditions in Theorem 1 (i.e., it can be shown that $\mathcal{N}(s)$ is not passive [24, 25]), we resort to dynamic loop-shifting techniques as presented in Section 2.3.
 More specifically, we consider the loop-shifted system $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$ in Fig.
-8(b) with the subsystems $\mathcal{D}^{\prime}(s)$ and $\mathcal{N}^{\prime}(s)$ given as
+9(b) with the subsystems $\mathcal{D}^{\prime}(s)$ and $\mathcal{N}^{\prime}(s)$ given as
 $\begin{split}\mathcal{D}^{\prime}(s)&\coloneqq\mathcal{D}(s)(I-\Gamma(s)\mathcal{D}(s))^{-1}\\ \mathcal{N}^{\prime}(s)&\coloneqq\mathcal{N}(s)+\Gamma(s).\end{split}$ (29)
-We choose $\Gamma(s)=\mathrm{diag}(\Gamma_{1}(s),...,\Gamma_{n}(s))$ as a block-diagonal semi-stable transfer matrix with the $2\times 2$ matrix elements
+Note again that loop-shifting with $\Gamma(s)$ is a mathematical tool for establishing passivity properties and does not involve any hardware or impedance modifications of the actual network or converter system.
+Having said that, we choose $\Gamma(s)=\mathrm{diag}(\Gamma_{1}(s),...,\Gamma_{n}(s))$ as a block-diagonal semi-stable transfer matrix with the $2\times 2$ matrix elements
 $\Gamma_{i}(s)=\begin{bmatrix}\Gamma_{i}^{\mathrm{p}}(s)&0\\ 0&\Gamma_{i}^{\mathrm{q}}(s)\end{bmatrix}$, (30)
@@
 $\begin{split}\Gamma_{i}^{\mathrm{p}}(s)&=\tfrac{1}{s}\left(\tfrac{\gamma_{1,i}^{\mathrm{p}}\omega_{0}^{2}s^{2}}{\omega_{0}^{2}+(\rho\omega_{0}+s)^{2}}+\tfrac{\gamma_{2,i}^{\mathrm{p}}\omega_{0}^{2}}{\omega_{0}^{2}+(\rho\omega_{0}+s)^{2}}+\gamma_{3,i}^{\mathrm{p}}\right)\\ \Gamma_{i}^{\mathrm{q}}(s)&=\tfrac{1}{s}\left(\tfrac{\gamma_{1,i}^{\mathrm{q}}\omega_{0}^{2}s^{2}}{\omega_{0}^{2}+(\rho\omega_{0}+s)^{2}}+\tfrac{\gamma_{2,i}^{\mathrm{q}}\omega_{0}^{2}}{\omega_{0}^{2}+(\rho\omega_{0}+s)^{2}}+\gamma_{3,i}^{\mathrm{q}}\right)\end{split}$ (31)
-with the parameters
-$\begin{split}\gamma_{1,i}^{\mathrm{p}}&=\gamma_{1,i}^{\mathrm{q}}=\tfrac{2|v|_{\mathrm{max}}^{2}}{\omega_{0}^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}\\ \gamma_{2,i}^{\mathrm{p}}&=\gamma_{2,i}^{\mathrm{q}}=-3|v|_{\mathrm{max}}^{2}\textstyle\sum_{j\neq i}^{n}b_{ij}\\ \gamma_{3,i}^{\mathrm{p}}&=3\tfrac{1}{1+\rho^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}|v|_{\mathrm{max}}^{2}\\ \gamma_{3,i}^{\mathrm{q}}&=\tfrac{1}{1+\rho^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}\left(3|v|_{\mathrm{max}}^{2}+0.8\right),\end{split}$ (31)
-whose structure loosely resembles the network model in 17.
+where the parameters are selected as
+$\begin{split}\gamma_{1,i}^{\mathrm{p}}&=\gamma_{1,i}^{\mathrm{q}}=\tfrac{2|v|_{\mathrm{max}}^{2}}{\omega_{0}^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}\\ \gamma_{2,i}^{\mathrm{p}}&=\gamma_{2,i}^{\mathrm{q}}=-3|v|_{\mathrm{max}}^{2}\textstyle\sum_{j\neq i}^{n}b_{ij}\\ \gamma_{3,i}^{\mathrm{p}}&=-\tfrac{\gamma_{2,i}^{\mathrm{p}}}{1+\rho^{2}}\\ \gamma_{3,i}^{\mathrm{q}}&=-\tfrac{\gamma_{2,i}^{\mathrm{q}}}{1+\rho^{2}}+\tilde{\gamma}_{3,i}^{\mathrm{q}},\end{split}$ (32)
+where $\tilde{\gamma}_{3,i}^{\mathrm{q}}\geq\tfrac{0.8}{1+\rho^{2}}\textstyle\sum_{j\neq i}^{n}b_{ij}$.
+The choice of the transfer function $\Gamma(s)$ is guided by structural insights into the dynamic network model in 18, and algebraic considerations to ensure passivity of the loop-shifted system $\mathcal{N}^{\prime}(s)$, as detailed below.
+[…]
+Figure 8: Structural overview of the proof of Theorem 2.
 III. Passivity Checks & Theorem 1
-For $\Gamma(s)$ selected as in 29, 30 and 31, we can show that $\mathcal{N}^{\prime}(s)$ is passive [^2] Notice that we can also show passivity of $\mathcal{N}^{\prime}(s)$ for the lossless case $\rho=0$ by evaluating 29, 30 and 31 at $\rho=0$ accordingly., i.e., it satisfies the conditions (i) to (iii) in Definition 2:
+For $\Gamma(s)$ selected as in 30, 31 and 32, we can show that $\mathcal{N}^{\prime}(s)$ is passive, i.e., it satisfies the conditions (i) to (iii) in Definition 2:
+(i) Poles: The poles of all elements of $\mathcal{N}^{\prime}(s)$ are $p_{1}=\mathrm{j}0$ and $p_{2,3}=-\rho\omega_{0}\pm\mathrm{j}\omega_{0}$, i.e., $\text{Re}(p_{k})\leq 0$ for $k\in\{1,2,3\}$.
+(ii) Positive semi-definiteness: We can express the Hermitian matrix $\mathcal{S}_{\mathcal{N}^{\prime}}(\mathrm{j}\omega)\coloneqq\mathcal{N}^{\prime}(\mathrm{j}\omega)+\mathcal{N}^{\prime\star}(\mathrm{j}\omega)$ as
 […]
-(ii) Positive semi-definiteness: We can express the Hermitian matrix $\mathcal{S}_{\mathcal{N}^{\prime}}(\mathrm{j}\omega)\coloneqq\mathcal{N}^{\prime}(\mathrm{j}\omega)+\mathcal{N}^{\prime\star}(\mathrm{j}\omega)$ as
-$\begin{split}\mathcal{S}_{\mathcal{N}^{\prime}}(\mathrm{j}\omega)=\begin{bmatrix}\mathcal{S}_{\mathcal{N}^{\prime},11}(\mathrm{j}\omega)&\dots&\mathcal{S}_{\mathcal{N}^{\prime},1n}(\mathrm{j}\omega)\\ \vdots&\ddots&\vdots\\ \mathcal{S}_{\mathcal{N}^{\prime},n1}(\mathrm{j}\omega)&\dots&\mathcal{S}_{\mathcal{N}^{\prime},nn}(\mathrm{j}\omega)\end{bmatrix},\end{split}$ (32)
 where each $\mathcal{S}_{\mathcal{N}^{\prime},ij}$ represents a $2\times 2$ transfer matrix block. The diagonal and off-diagonal elements are given by
-$\begin{split}\mathcal{S}_{\mathcal{N}^{\prime},ii}(\mathrm{j}\omega)&=h_{\rho}(\omega)\left(\textstyle\sum_{j\neq i}^{n}|v|_{0,i}^{2}b_{ij}\begin{bmatrix}-1&-\mathrm{j}\omega/\omega_{0}\\ \mathrm{j}\omega/\omega_{0}&-1\end{bmatrix}+\right.\\ &\quad\quad\quad\quad\quad\,\,\,\left.\begin{bmatrix}\omega^{2}\gamma_{1,i}^{\mathrm{p}}-\gamma_{2,i}^{\mathrm{p}}&0\\ 0&\omega^{2}\gamma_{1,i}^{\mathrm{q}}-\gamma_{2,i}^{\mathrm{q}}\end{bmatrix}\right)\\ \mathcal{S}_{\mathcal{N}^{\prime},ij}(\mathrm{j}\omega)&=h_{\rho}(\omega)|v|_{0,i}|v|_{0,j}b_{ij}\begin{bmatrix}1&\mathrm{j}\omega/\omega_{0}\\ -\mathrm{j}\omega/\omega_{0}&1\end{bmatrix}.\end{split}$ (33)
+[…]
 Here, $h_{\rho}(\omega)$ is a positive function $\forall\omega\geq 0$ defined as
-$h_{\rho}(\omega)={4\rho\omega_{0}^{3}}/({(\omega_{0}^{2}+\rho^{2}\omega_{0}^{2}-\omega^{2})^{2}+4\rho^{2}\omega_{0}^{2}\omega^{2}})\geq 0.$ (34)
-We can observe that 32 is a Hermitian diagonally dominant matrix with real non-negative diagonal entries, i.e., the magnitude of the diagonal entry in a row is greater or equal to the sum of the magnitudes of the off-diagonal entries in that row.
+$\hskip-5.69054pth_{\rho}(\omega)={4\rho\omega_{0}^{3}}/({(\omega_{0}^{2}\hskip-1.42262pt+\hskip-1.42262pt\rho^{2}\omega_{0}^{2}\hskip-1.42262pt-\hskip-1.42262pt\omega^{2})^{2}\hskip-1.42262pt+\hskip-1.42262pt4\rho^{2}\omega_{0}^{2}\omega^{2}})\geq 0.$ (35)
+We can observe that 33 is a Hermitian diagonally dominant matrix with real non-negative diagonal entries, i.e., the magnitude of the diagonal entry in a row is greater or equal to the sum of the magnitudes of the off-diagonal entries in that row.
 Namely, for the odd rows and $\forall\omega\geq 0$, we get
-$\begin{split}|\textstyle\sum_{j\neq i}^{n}-|v|_{0,i}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{p}}-\gamma_{2,i}^{\mathrm{p}}|&\geq\\ |\textstyle\sum_{j\neq i}^{n}|v|_{0,i}^{2}b_{ij}\tfrac{-\mathrm{j}\omega}{\omega_{0}}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}&|v|_{0,j}b_{ij}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}|v|_{0,j}b_{ij}\tfrac{\mathrm{j}\omega}{\omega_{0}}|\\ \Leftrightarrow|\textstyle\sum_{j\neq i}^{n}-|v|_{\mathrm{max}}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{p}}-\gamma_{2,i}^{\mathrm{p}}|&\geq\\ 2\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}&b_{ij}\tfrac{\omega}{\omega_{0}}+\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}\hskip 42.67912pt\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(2+\tfrac{2\omega^{2}}{\omega_{0}^{2}})&\geq\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(1+\tfrac{2\omega}{\omega_{0}}),\end{split}$
+$\begin{split}|\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt-|v|_{0,i}^{2}b_{ij}\hskip-1.42262pt+\hskip-1.42262pt\omega^{2}\gamma_{1,i}^{\mathrm{p}}\hskip-1.42262pt-\hskip-1.42262pt\gamma_{2,i}^{\mathrm{p}}|\hskip-1.70717pt&\geq\\ |\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{0,i}^{2}b_{ij}\tfrac{-\mathrm{j}\omega}{\omega_{0}}|\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt|\hskip-0.85358pt|v|_{0,i}&|v|_{0,j}b_{ij}|\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|\hskip-0.85358pt|v|_{0,i}|v|_{0,j}b_{ij}\tfrac{\mathrm{j}\omega}{\omega_{0}}|\hskip-5.69054pt\\ \Leftrightarrow\hskip-1.42262pt|\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt-|v|_{\mathrm{max}}^{2}b_{ij}\hskip-1.42262pt+\hskip-1.42262pt\omega^{2}\gamma_{1,i}^{\mathrm{p}}\hskip-1.42262pt-\hskip-1.42262pt\gamma_{2,i}^{\mathrm{p}}|\hskip-1.70717pt&\geq\hskip-1.70717pt\\ 2\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt\hskip-1.42262pt|v|_{\mathrm{max}}^{2}&b_{ij}\tfrac{\omega}{\omega_{0}}\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt\hskip-1.42262pt|v|_{\mathrm{max}}^{2}b_{ij}\hskip 42.67912pt\\ \Leftrightarrow\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}b_{ij}(2+\tfrac{2\omega^{2}}{\omega_{0}^{2}})\hskip-1.70717pt&\geq\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}b_{ij}\hskip-0.56905pt(1+\tfrac{2\omega}{\omega_{0}}),\end{split}$
 and similarly, for each even row and $\forall\omega\geq 0$, we have
-$\begin{split}|\textstyle\sum_{j\neq i}^{n}-|v|_{0,i}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{q}}-\gamma_{2,i}^{\mathrm{q}}|&\geq\\ |\textstyle\sum_{j\neq i}^{n}|v|_{0,i}^{2}b_{ij}\tfrac{\mathrm{j}\omega}{\omega_{0}}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}&|v|_{0,j}b_{ij}\tfrac{-\mathrm{j}\omega}{\omega_{0}}|+\textstyle\sum_{j\neq i}^{n}||v|_{0,i}|v|_{0,j}b_{ij}|\\ \Leftrightarrow|\textstyle\sum_{j\neq i}^{n}-|v|_{\mathrm{max}}^{2}b_{ij}+\omega^{2}\gamma_{1,i}^{\mathrm{q}}-\gamma_{2,i}^{\mathrm{q}}|&\geq\\ 2\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}&b_{ij}\tfrac{\omega}{\omega_{0}}+\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(2+\tfrac{2\omega^{2}}{\omega_{0}^{2}})&\geq\textstyle\sum_{j\neq i}^{n}|v|_{\mathrm{max}}^{2}b_{ij}(1+\tfrac{2\omega}{\omega_{0}}).\end{split}$
+$\begin{split}\hskip-5.69054pt|\hskip-0.85358pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt-|v|_{0,i}^{2}b_{ij}\hskip-1.42262pt+\hskip-1.42262pt\omega^{2}\gamma_{1,i}^{\mathrm{q}}\hskip-1.42262pt-\hskip-1.42262pt\gamma_{2,i}^{\mathrm{q}}|&\hskip-1.70717pt\geq\hskip-1.70717pt\\ |\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{0,i}^{2}b_{ij}\tfrac{\mathrm{j}\omega}{\omega_{0}}|\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt|\hskip-0.85358pt|v|_{0,i}&|v|_{0,j}b_{ij}\tfrac{-\mathrm{j}\omega}{\omega_{0}}\hskip-0.85358pt|\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|\hskip-0.85358pt|v|_{0,i}|v|_{0,j}b_{ij}\hskip-0.85358pt|\hskip-1.99168pt\hskip-5.69054pt\\ \hskip-5.69054pt\Leftrightarrow|\hskip-0.85358pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt-|v|_{\mathrm{max}}^{2}b_{ij}\hskip-1.42262pt+\hskip-1.42262pt\omega^{2}\gamma_{1,i}^{\mathrm{q}}\hskip-1.42262pt-\hskip-1.42262pt\gamma_{2,i}^{\mathrm{q}}|&\hskip-1.70717pt\geq\hskip-1.70717pt\\ 2\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}&b_{ij}\tfrac{\omega}{\omega_{0}}\hskip-1.99168pt+\hskip-1.99168pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}b_{ij}\\ \hskip-5.69054pt\Leftrightarrow\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}b_{ij}(2+\tfrac{2\omega^{2}}{\omega_{0}^{2}})&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt|v|_{\mathrm{max}}^{2}b_{ij}\hskip-0.56905pt(1+\tfrac{2\omega}{\omega_{0}}).\end{split}$
 By Lemma 1, we can conclude $\mathcal{N}^{\prime}(\mathrm{j}\omega)+\mathcal{N}^{\prime\star}(\mathrm{j}\omega)\succeq 0$.
-$\mathcal{D}(s)$$\mathcal{N}(s)$--$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|\tilde{v}|_{\mathrm{n}}\end{bmatrix}$$\begin{bmatrix}\Delta p\\ \Delta q\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|\tilde{v}|_{\mathrm{nd}}\end{bmatrix}$device dynamicsnetwork dynamics--$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$
+(iii) Imaginary poles: For $\rho\neq 0$, $\mathcal{N}^{\prime}(s)$ has one imaginary pole, i.e., $p_{1}=\mathrm{j}0$, which is a simple pole.
+We therefore compute the limit $\mathcal{R}_{\mathrm{j}0}^{\mathcal{N}^{\prime}}\coloneqq\text{lim}_{s\rightarrow\mathrm{j}0}(s-\mathrm{j}0)\mathcal{N}^{\prime}(s)$, where each $\mathcal{R}_{\mathrm{j}0,ij}^{\mathcal{N}^{\prime}}$ represents a $2\times 2$ transfer matrix block.
+The diagonal and off-diagonal elements are given by
+[…]
+Again, 36 is a Hermitian diagonally dominant matrix with real non-negative diagonal entries. For the odd rows, we get
+$\begin{split}\hskip-5.69054pt|\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717ptb_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\hskip-1.42262pt+\hskip-1.42262pt\tfrac{\gamma_{2,i}^{\mathrm{p}}}{1+\rho^{2}}\hskip-1.42262pt+\hskip-1.42262pt\gamma_{3,i}^{\mathrm{p}}|&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt|\hskip-0.85358pt-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\hskip-1.99168pt\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}\hskip-1.70717ptb_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717ptb_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}.\end{split}$
+For the even rows we get (with $|v|_{\mathrm{max}}\hskip-1.42262pt=\hskip-1.42262pt1.1$ and $|v|_{\mathrm{min}}\hskip-1.42262pt=\hskip-1.42262pt0.9$):
+$\begin{split}\hskip-5.69054pt|\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt\tfrac{(2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j})b_{ij}}{1+\rho^{2}}\hskip-1.42262pt+\hskip-1.42262pt\tfrac{\gamma_{2,i}^{\mathrm{q}}}{1+\rho^{2}}\hskip-1.42262pt+\hskip-1.42262pt\gamma_{3,i}^{\mathrm{q}}|&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt|b_{ij}\tfrac{-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\hskip-1.99168pt\\ \Leftrightarrow|\hskip-1.42262pt\textstyle\sum_{j\neq i}^{n}\hskip-1.70717pt\tfrac{(2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j})b_{ij}}{1+\rho^{2}}+\tilde{\gamma}_{3,i}^{\mathrm{q}}|&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt|b_{ij}\tfrac{-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\hskip-1.99168pt\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}\hskip-1.70717ptb_{ij}\tfrac{2|v|_{\mathrm{min}}^{2}-|v|_{\mathrm{max}}^{2}}{1+\rho^{2}}+\tilde{\gamma}_{3,i}^{\mathrm{q}}&\hskip-1.70717pt\geq\hskip-1.70717pt\textstyle\sum_{j\neq i}^{n}\hskip-0.85358pt\hskip-0.85358ptb_{ij}\tfrac{|v|_{\mathrm{max}^{2}}}{1+\rho^{2}}\hskip-1.99168pt\end{split}$
+By Lemma 1, we conclude $\mathcal{R}_{\mathrm{j}0}^{\mathcal{N}^{\prime}}\succeq 0$.
+$\mathcal{D}(s)$$\mathcal{N}(s)$-$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|\tilde{v}|_{\mathrm{n}}\end{bmatrix}$$\begin{bmatrix}\Delta p\\ \Delta q\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|\tilde{v}|_{\mathrm{nd}}\end{bmatrix}$device dynamicsnetwork dynamics-$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$
 (a) Original feedback system.
-$\mathcal{D}(s)$$\mathcal{D}^{\prime}$$\mathcal{N}(s)$$\mathcal{N}^{\prime}$--$\Gamma(s)$$\Gamma(s)$$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|\tilde{v}|_{\mathrm{n}}\end{bmatrix}$device dynamicsnetwork dynamics$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|\tilde{v}|_{\mathrm{nd}}\end{bmatrix}$
+$\mathcal{D}(s)$$\mathcal{D}^{\prime}$$\mathcal{N}(s)$$\mathcal{N}^{\prime}$-$\Gamma(s)$$\Gamma(s)$$\begin{bmatrix}\Delta p_{\mathrm{d}}\\ \Delta q_{\mathrm{d}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega\\ \Delta|\tilde{v}|_{\mathrm{n}}\end{bmatrix}$device dynamicsnetwork dynamics[$\begin{bmatrix}\Delta p_{\mathrm{e}}\\ \Delta q_{\mathrm{e}}\end{bmatrix}$$\begin{bmatrix}\Delta\omega_{\mathrm{d}}\\ \Delta|\tilde{v}|_{\mathrm{nd}}\end{bmatrix}$
 (b) Loop-shifting with $\Gamma$.
 Figure 9: Closed-loop feedback interconnection for stability analysis.
+Next, we derive the decentralized stability conditions in 25 and 26 under which $\mathcal{D}^{\prime}(s)$ is strictly passive.
+We consider $\mathcal{D}^{\prime}(s)=\text{diag}(\mathcal{D}_{1}^{\prime}(s),\dots,\mathcal{D}_{n}^{\prime}(s))$ with the matrix elements
 […]
-$\begin{split}\mathcal{R}_{\mathrm{j}0,ii}^{\mathcal{N}^{\prime}}&=\textstyle\sum_{j\neq i}^{n}b_{ij}\begin{bmatrix}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}&0\\ 0&\tfrac{2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\end{bmatrix}+\\ &\quad\quad\quad\quad\quad\quad\quad\quad\,\,\,\begin{bmatrix}\tfrac{\gamma_{2,i}^{\mathrm{p}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{p}}&0\\ 0&\tfrac{\gamma_{2,i}^{\mathrm{q}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{q}}\end{bmatrix}\\ \mathcal{R}_{\mathrm{j}0,ij}^{\mathcal{N}^{\prime}}&=-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}\begin{bmatrix}1&0\\ 0&1\end{bmatrix}.\end{split}$ (35)
-Again, 35 is a Hermitian diagonally dominant matrix with real non-negative diagonal entries. For the odd rows, we get
-$\begin{split}|\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}+\tfrac{\gamma_{2,i}^{\mathrm{p}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{p}}|&\geq\textstyle\sum_{j\neq i}^{n}|-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}&\geq\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}.\end{split}$
-For the even rows we get (with $|v|_{\mathrm{max}}=1.1$ and $|v|_{\mathrm{min}}=0.9$):
-$\begin{split}|\textstyle\sum_{j\neq i}^{n}\tfrac{(2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j})b_{ij}}{1+\rho^{2}}+\tfrac{\gamma_{2,i}^{\mathrm{q}}}{1+\rho^{2}}+\gamma_{3,i}^{\mathrm{q}}|&\geq\textstyle\sum_{j\neq i}^{n}|b_{ij}\tfrac{-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\\ \Leftrightarrow|\textstyle\sum_{j\neq i}^{n}\tfrac{(2|v|_{0,i}^{2}-|v|_{0,i}|v|_{0,j}+0.8)b_{ij}}{1+\rho^{2}}|&\geq\textstyle\sum_{j\neq i}^{n}|b_{ij}\tfrac{-|v|_{0,i}|v|_{0,j}}{1+\rho^{2}}|\\ \Leftrightarrow\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{2|v|_{\mathrm{min}}^{2}-|v|_{\mathrm{max}}^{2}+0.8}{1+\rho^{2}}&\geq\textstyle\sum_{j\neq i}^{n}b_{ij}\tfrac{|v|_{\mathrm{max}^{2}}}{1+\rho^{2}}\end{split}$
-By Lemma 1, we conclude $\mathcal{R}_{\mathrm{j}0}^{\mathcal{N}^{\prime}}\succeq 0$.
-Next, we derive the decentralized stability conditions in 24 and 25 under which $\mathcal{D}^{\prime}(s)$ is strictly passive.
-We consider $\mathcal{D}^{\prime}(s)=\text{diag}(\mathcal{D}_{1}^{\prime}(s),\dots,\mathcal{D}_{n}^{\prime}(s))$ with the matrix elements
-$\begin{split}\mathcal{D}_{i}^{\prime}(s)=\mathcal{D}_{i}(s)(I-\Gamma_{i}(s)\mathcal{D}_{i}(s))^{-1}=\begin{bmatrix}\mathcal{D}_{\mathrm{p},i}^{\prime}(s)&0\\ 0&\mathcal{D}^{\prime}_{\mathrm{q},i}(s)\end{bmatrix},\end{split}$ (36)
 where the diagonal transfer function elements are given as
 $\begin{split}\mathcal{D}_{\mathrm{p},i}^{\prime}(s)&=\tfrac{d_{\mathrm{p},i}}{\tau_{\mathrm{p},i}s+1-\Gamma_{i}^{\mathrm{p}}(s)d_{\mathrm{p},i}}\end{split}$ (38a)
-$\begin{split}\mathcal{D}_{\mathrm{q},i}^{\prime}(s)&=\tfrac{d_{\mathrm{q},i}s/|v|_{0,i}}{\tau_{\mathrm{q},i}s+1-\Gamma_{i}^{\mathrm{q}}(s)d_{\mathrm{q},i}s/|v|_{0,i}},\end{split}$ (37b)
-for each of which we check the strict passivity conditions (i) and (ii) in Definition 3 independently. We start with $\mathcal{D}_{\mathrm{p},i}^{\prime}(s)$:
-(i) Poles: To show that the poles of all elements of $\mathcal{D}_{\mathrm{p},i}^{\prime}(s)$ are in $\text{Re}(s)<0$, we start by inserting the expression for $\Gamma_{i}^{\mathrm{p}}(s)$ in 30 into 37a and obtain
+$\begin{split}\mathcal{D}_{\mathrm{q},i}^{\prime}(s)&=\tfrac{\tfrac{d_{\mathrm{q},i}s}{|v|_{0,i}}}{\tau_{\mathrm{q},i}s+1-\Gamma_{i}^{\mathrm{q}}(s)\tfrac{d_{\mathrm{q},i}s}{|v|_{0,i}}},\end{split}$ (38b)
+for each of which we check the strict passivity conditions (i) and (ii) in Definition 3 independently.
+We start with $\mathcal{D}_{\mathrm{p},i}^{\prime}(s)$ and insert the expression for $\Gamma_{i}^{\mathrm{p}}(s)$ in 31 into 38a, i.e.,
 $\mathcal{D}_{\mathrm{p},i}^{\prime}(s)=d_{\mathrm{p},i}\tfrac{s^{2}b_{2,i}+sb_{1,i}+b_{0,i}}{s^{3}a_{3,i}+s^{2}a_{2,i}+sa_{1,i}+a_{0,i}},$ (39)
 where the transfer function coefficients are given by
-$\begin{split}a_{0,i}&=\omega_{0}^{2}(\rho^{4}+2\rho^{2}-6\rho\alpha_{\mathrm{p},i}+1)\\ a_{1,i}&=\omega_{0}(\tilde{\tau}_{\mathrm{p},i}\rho^{4}+2\rho^{3}+2\rho^{2}\tilde{\tau}_{\mathrm{p},i}\\ &-2\rho^{2}\alpha_{\mathrm{p},i}+2\rho-5\alpha_{\mathrm{p},i}+\tilde{\tau}_{\mathrm{p},i})\\ a_{2,i}&=(1+\rho^{2})(2\tilde{\tau}_{\mathrm{p},i}\rho+1)\\ a_{3,i}&=\tau_{\mathrm{p},i}(1+\rho^{2})\end{split}\begin{split}b_{0,i}&=\omega_{0}^{2}(1+\rho^{2})^{2}\\ b_{1,i}&=2\rho\omega_{0}(1+\rho^{2})\\ b_{2,i}&=1+\rho^{2}.\end{split}$ (39)
-By checking the Hurwitz criterion for a 3rd order polynomial, we can conclude that $a_{3,i}>0$ and $a_{2,i}>0$ are always satisfied, while $a_{1,i}>0$ is ensured by condition 24a, $a_{0,i}>0$ by condition 24b, and $a_{2,i}a_{1,i}>a_{0,i}a_{3,i}$ by condition 24c.
-(ii) Positive definiteness: To ensure $\mathcal{D}_{\mathrm{p},i}^{\prime}(\mathrm{j}\omega)+\mathcal{D}^{\prime\star}_{\mathrm{p},i}(\mathrm{j}\omega)\succ 0$, we require the additional condition 24d.
-Following the same reasoning, we derive the decentralized stability conditions [^3] The requirement $\tau_{\mathrm{q},i}\neq 0$ in 25 ensures properness of $\mathcal{D}_{\mathrm{q},i}^{\prime}(s)$.
-for the reactive power-voltage droop control presented in 25 by checking strict passivity of $\mathcal{D}_{\mathrm{q},i}^{\prime}(s)$.
-Finally, since $\mathcal{N}^{\prime}(s)$ is passive and $\mathcal{D}^{\prime}(s)$ is strictly passive, we can apply Theorem 1 and conclude $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}\in\mathcal{RH}_{\infty}^{4n\times 4n}$.
+$\begin{split}\hskip-2.84526pta_{0,i}&\hskip-1.42262pt=\hskip-1.42262pt\omega_{0}^{2}(1\hskip-1.42262pt+\hskip-1.42262pt\rho^{2})-2\rho\omega_{0}d_{\mathrm{p},i}\gamma_{3,i}^{\mathrm{p}}\\ \hskip-2.84526pta_{1,i}&\hskip-1.42262pt=2\rho\omega_{0}+\tau_{\mathrm{p,i}}\omega_{0}^{2}(1\hskip-1.42262pt+\hskip-1.42262pt\rho^{2})\hskip-1.42262pt\\ &\hskip-1.42262pt-d_{\mathrm{p},i}\omega_{0}^{2}\gamma_{1,i}^{\mathrm{p}}-d_{\mathrm{p},i}\gamma_{3,i}^{\mathrm{p}}\\ \hskip-2.84526pta_{2,i}&\hskip-1.42262pt=\hskip-1.42262pt2\omega_{0}{\tau}_{\mathrm{p},i}\rho\hskip-1.42262pt+\hskip-1.42262pt1\\ \hskip-2.84526pta_{3,i}&\hskip-1.42262pt=\hskip-1.42262pt\tau_{\mathrm{p},i}\end{split}\begin{split}b_{0,i}&\hskip-1.42262pt=\hskip-1.42262pt\omega_{0}^{2}(1\hskip-1.42262pt+\hskip-1.42262pt\rho^{2})\hskip-2.84526pt\\ b_{1,i}&\hskip-1.42262pt=\hskip-1.42262pt2\rho\omega_{0}\hskip-2.84526pt\\ b_{2,i}&\hskip-1.42262pt=\hskip-1.42262pt1.\end{split}$ (40)
+(i) Poles: To show that the poles of all elements of $\mathcal{D}_{\mathrm{p},i}^{\prime}(s)$ are in $\text{Re}(s)<0$, we check the Hurwitz criterion for a 3rd order polynomial, and conclude that $a_{3,i}>0$ and $a_{2,i}>0$ are always satisfied, while $a_{1,i}>0$ is ensured by 25a, $a_{0,i}>0$ by 25b, and $a_{2,i}a_{1,i}>a_{0,i}a_{3,i}$ by 25c.
+(ii) Positive definiteness: To ensure $\mathcal{D}_{\mathrm{p},i}^{\prime}(\mathrm{j}\omega)+\mathcal{D}^{\prime\star}_{\mathrm{p},i}(\mathrm{j}\omega)\succ 0,\,\forall\omega\in(-\infty,\infty)$, we require the additional condition 25d.
+Following the same reasoning, we derive the decentralized stability conditions for the reactive power-voltage droop control in 26 by checking strict passivity of $\mathcal{D}_{\mathrm{q},i}^{\prime}(s)$.
+To do so, we select $\tfrac{d_{\mathrm{q},i}}{|v|_{0,i}}=\tfrac{1}{\tilde{\gamma}_{3,i}^{\mathrm{q}}}$ to cancel the zero at the origin, and ensure strict passivity of the minimal realization of $\mathcal{D}_{\mathrm{q},i}^{\prime}(s)$.
+Finally, since $\mathcal{N}^{\prime}(s)$ is passive and $\mathcal{D}^{\prime}(s)$ is strictly passive, and additionally $\bar{\sigma}(\mathcal{N}^{\prime}(\mathrm{j}\infty))\bar{\sigma}(\mathcal{D}^{\prime}(\mathrm{j}\infty))<1$ (because $\mathcal{N}^{\prime}(\mathrm{j}\infty)=0_{2n\times 2n}$), we can apply Theorem 1 and conclude $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}\in\mathcal{RH}_{\infty}^{4n\times 4n}$.
 IV. Final Value Theorem & Lemma 2
-The internal feedback stability of the original system $\mathcal{D}_{0}\#\mathcal{N}_{0}$ follows directly from the internal feedback stability of the loop-shifted system $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$, as stated in the following lemma.
+The internal feedback stability of the original system $\mathcal{D}_{0}\#\mathcal{N}_{0}$ follows directly from the internal feedback stability of the minimal realization of all four closed-loop transfer functions of $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$, as stated in the following lemma.
 Lemma 4.
-Consider the device and network dynamics $\mathcal{D}^{\prime}(s)$ and $\mathcal{N}^{\prime}(s)$ in 28, as well as $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$ in 23 with the network model 17.
-Then, internal feedback stability of $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$ implies internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$.
+Consider the device and network dynamics $\mathcal{D}^{\prime}(s)$ and $\mathcal{N}^{\prime}(s)$ in 29, as well as $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$ in 24 with the network model 18.
+Then, internal feedback stability of the minimal realization of all four closed-loop transfer functions of the loop-shifted system $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$ implies internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$.
 Proof.
-We first conclude that $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}\in\mathcal{RH}_{\infty}^{4n\times 4n}$ implies $(I+\mathcal{D}^{\prime}\mathcal{N}^{\prime})^{-1}\mathcal{D}^{\prime}=(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}\in\mathcal{RH}_{\infty}^{2n\times 2n}$.
+We first conclude that $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}\in\mathcal{RH}_{\infty}^{4n\times 4n}$ implies $(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}^{\prime}\mathcal{N}^{\prime})^{-1}\mathcal{D}^{\prime}\hskip-1.42262pt=\hskip-1.42262pt(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}\mathcal{N})^{-1}\mathcal{D}\hskip-1.42262pt\in\hskip-1.42262pt\mathcal{RH}_{\infty}^{2n\times 2n}$, where the latter closed-loop transfer functions are minimal realizations.
 Further, since
 $\begin{split}\mathcal{D}(s)&=\mathcal{D}_{0}(s)\cdot\text{diag}(1,\tfrac{s}{|v|_{0,1}},\dots,1,\tfrac{s}{|v|_{0,n}})\\ \mathcal{N}(s)&=\mathcal{N}_{0}(s)\cdot\text{diag}(1,\tfrac{|v|_{0,1}}{s},\dots,1,\tfrac{|v|_{0,n}}{s}),\end{split}$ (41)
-$(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}=\text{diag}(1,\tfrac{s}{|v|_{0,1}},\dots,1,\tfrac{s}{|v|_{0,n}})(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}$, and we can thus conclude stability of $(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}$ if the step response of the voltage derivatives converges to zero, i.e.,
+$(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}\mathcal{N})^{-1}\mathcal{D}\hskip-1.42262pt=\hskip-1.42262pt\text{diag}(1,\tfrac{s}{|v|_{0,1}},\dots,1,\tfrac{s}{|v|_{0,n}})(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}$, and we can thus conclude stability of $(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}$ if the step response of the voltage derivatives converges to zero, i.e.,
 $\underset{s\rightarrow 0}{\text{lim}}\,\,s\frac{1}{s}(I+\mathcal{D}(s)\mathcal{N}(s))^{-1}\mathcal{D}(s)=I\otimes\begin{bmatrix}\star&0\\ 0&0\end{bmatrix}$. (42)
-To show that 41 holds, we first rewrite the transfer matrix $(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}=(I+\mathcal{D}{N}\tfrac{1}{s})^{-1}\mathcal{D}$, and then use the row permutation matrix $\mathcal{P}\in\mathbb{R}^{2n\times 2n}$ with elements
+To show that 42 holds, we first rewrite the transfer matrix $(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}=(I+\mathcal{D}{N}\tfrac{1}{s})^{-1}\mathcal{D}$, and then use the row permutation matrix $\mathcal{P}\in\mathbb{R}^{2n\times 2n}$ with elements
 $\mathcal{P}_{ij}=\begin{cases}1\quad\quad i=k,\,j=2k-1,\,1\leq k\leq n\\ 1\quad\quad i=k+n,\,j=2k,\,1\leq k\leq n\\ 0\quad\quad\text{else},\end{cases}$ (43)
@@
 The row-permutated matrices $\mathcal{D}^{\mathcal{P}}(s)$ and $N^{\mathcal{P}}(s)$ are given as
-$\mathcal{D}^{\mathcal{P}}(s)=\begin{bmatrix}\mathcal{D}_{\mathrm{p}}^{\mathcal{P}}(s)&0_{n\times n}\\ 0_{n\times n}&\mathcal{D}_{\mathrm{q}}^{\mathcal{P}}(s)\end{bmatrix},\,\,N^{\mathcal{P}}(s)=\begin{bmatrix}N_{1}^{\mathcal{P}}(s)&N_{2}^{\mathcal{P}}(s)\\ -N_{2}^{\mathcal{P}}(s)&N_{3}^{\mathcal{P}}(s)\end{bmatrix}$ (43)
+$\hskip-2.84526pt\mathcal{D}^{\mathcal{P}}\hskip-1.42262pt(s)\hskip-1.42262pt=\hskip-2.84526pt\begin{bmatrix}\mathcal{D}_{\mathrm{p}}^{\mathcal{P}}\hskip-1.42262pt(s)\hskip-1.42262pt&0_{n\times n}\\ 0_{n\times n}&\hskip-1.42262pt\mathcal{D}_{\mathrm{q}}^{\mathcal{P}}\hskip-1.42262pt(s)\end{bmatrix}\hskip-2.84526pt,\,\,N^{\mathcal{P}}\hskip-1.42262pt(s)\hskip-1.42262pt=\hskip-2.84526pt\begin{bmatrix}N_{1}^{\mathcal{P}}(s)\hskip-1.42262pt&\hskip-1.42262ptN_{2}^{\mathcal{P}}\hskip-1.42262pt(s)\\ -N_{2}^{\mathcal{P}}\hskip-1.42262pt(s)\hskip-1.42262pt&\hskip-1.42262ptN_{3}^{\mathcal{P}}(s)\end{bmatrix}\hskip-2.84526pt$ (44)
 where $\mathcal{D}_{\mathrm{p}}^{\mathcal{P}}(s)=\text{diag}(\mathcal{D}_{\mathrm{p},1}(s),\dots,\mathcal{D}_{\mathrm{p},n}(s))$ and $\mathcal{D}_{\mathrm{q}}^{\mathcal{P}}(s)=\text{diag}(\mathcal{D}_{\mathrm{q},1}(s),\dots,\mathcal{D}_{\mathrm{q},n}(s))$ with elements
@@
 $N_{3,ij}^{\mathcal{P}}(s)$ $=-b_{ij}\tfrac{|v|_{0,i}|v|_{0,j}}{1+(\rho+\tfrac{s}{\omega_{0}})^{2}}.$
-Recalling that $(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}$ is stable, and thus its permuted version, we can apply the Final Value Theorem (FVT):
-$\begin{split}&\,\,\underset{s\rightarrow 0}{\text{lim}}\,\,(I+\mathcal{D}^{\mathcal{P}}(s){N}^{\mathcal{P}}(s)\tfrac{1}{s})^{-1}\mathcal{D}^{\mathcal{P}}(s)\\ =\,&\,\,\underset{s\rightarrow 0}{\text{lim}}\,\,(sI+\mathcal{D}^{\mathcal{P}}(s){N}^{\mathcal{P}}(s))^{-1}\mathcal{D}^{\mathcal{P}}(s)s\\ =\,&\,\,\underset{s\rightarrow 0}{\text{lim}}\,\left[\begin{array}[]{c}(sI+\mathcal{D}^{\mathcal{P}}_{\mathrm{p}}(s){N}^{\mathcal{P}}_{1}(s))^{-1}\mathcal{D}^{\mathcal{P}}_{\mathrm{p}}(s)s\\ 0_{n\times n}\end{array}\right.\\ \quad&\hskip 71.13188pt\left.\begin{array}[]{c}0_{n\times n}\\ (sI+\mathcal{D}^{\mathcal{P}}_{\mathrm{q}}(s){N}^{\mathcal{P}}_{3}(s))^{-1}\mathcal{D}^{\mathcal{P}}_{\mathrm{q}}(s)s\end{array}\right]\\ =\,&\begin{bmatrix}\star_{n\times n}&0_{n\times n}\\ 0_{n\times n}&0_{n\times n}\end{bmatrix},\end{split}$ (46)
-where for the second equality we have used $N_{2}^{\mathcal{P}}(0)=0_{n\times n}$.
-The last equality follows from the fact that $N_{1}^{\mathcal{P}}(0)$ is a Laplacian matrix with zero eigenvalue [22], while $N_{3}^{\mathcal{P}}(0)$ is a regular matrix which does not have a zero eigenvalue.
-Given 46, 41 follows, and thus $(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}\in\mathcal{RH}_{\infty}^{2n\times 2n}$.
+Recalling that $(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}\mathcal{N})^{-1}\mathcal{D}$ is stable, and thus its permuted version, we can apply the Final Value Theorem (FVT):
+$\,\,\underset{s\rightarrow 0}{\text{lim}}\,\,(I+\mathcal{D}^{\mathcal{P}}(s){N}^{\mathcal{P}}(s)\tfrac{1}{s})^{-1}\mathcal{D}^{\mathcal{P}}(s)$ (47a)
+$=\,$ $\,\,\underset{s\rightarrow 0}{\text{lim}}\,\,(sI+\mathcal{D}^{\mathcal{P}}(s){N}^{\mathcal{P}}(s))^{-1}\mathcal{D}^{\mathcal{P}}(s)s$ (47b)
+$=\,$ $\,\,\underset{s\rightarrow 0}{\text{lim}}\,\left[\begin{array}[]{c}\hskip-4.26773pt(sI+\mathcal{D}^{\mathcal{P}}_{\mathrm{p}}(s){N}^{\mathcal{P}}_{1}(s))^{-1}\mathcal{D}^{\mathcal{P}}_{\mathrm{p}}(s)s\\ \hskip-4.26773pt0_{n\times n}\end{array}\right$.
+(47e)
+$\hskip 71.13188pt\left.\begin{array}[]{c}0_{n\times n}\hskip-4.26773pt\\ (sI+\mathcal{D}^{\mathcal{P}}_{\mathrm{q}}(s){N}^{\mathcal{P}}_{3}(s))^{-1}\mathcal{D}^{\mathcal{P}}_{\mathrm{q}}(s)s\hskip-4.26773pt\end{array}\right]\hskip-2.84544pt$ (47h)
+$=\,$ $\begin{bmatrix}\star_{n\times n}&0_{n\times n}\\ 0_{n\times n}&0_{n\times n}\end{bmatrix}$, (47i)
+where for 47h we have used $N_{2}^{\mathcal{P}}(0)=0_{n\times n}$.
+The last equality in 47i follows from the fact that $N_{1}^{\mathcal{P}}(0)$ is a Laplacian matrix with zero eigenvalue [26], while $N_{3}^{\mathcal{P}}(0)$ is a regular matrix which does not have a zero eigenvalue.
+Given 47, 42 follows, and thus $(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}\in\mathcal{RH}_{\infty}^{2n\times 2n}$.
 Finally, since $\mathcal{D}_{0}(s)$ does not include a RHP pole or zero, there are no RHP pole-zero cancellations between $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$.
@@
 ∎
-(a) Stability conditions in 24.
-(b) Stability conditions in 25.
-Figure 9: 2D plot of the stability conditions for the three-node system with $\rho=0.05$. The stars indicate the feasibility of the GFM 3 control parameters.
+Remark 6.
+Recall again that we concluded stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}\in\mathcal{RH}_{\infty}^{4n\times 4n}$ by first proving stability of the minimal realization of all four closed-loop transfer functions of $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$, which implies stability of $(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}^{\prime}\mathcal{N}^{\prime})^{-1}\mathcal{D}^{\prime}\hskip-1.42262pt=\hskip-1.42262pt(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}\mathcal{N})^{-1}\mathcal{D}\hskip-1.42262pt\in\hskip-1.42262pt\mathcal{RH}_{\infty}^{2n\times 2n}$, where the latter closed-loop transfer functions are minimal realizations.
+Notice that there are RHP pole-zero cancellations between $\mathcal{D}(s)$ and $\mathcal{N}(s)$, i.e., we can NOT conclude stability of $\mathcal{D}\#\mathcal{N}$.
+However, from the stability of $\hskip-1.42262pt(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}\mathcal{N})^{-1}\mathcal{D}\hskip-1.42262pt\in\hskip-1.42262pt\mathcal{RH}_{\infty}^{2n\times 2n}$, we can conclude stability of $\hskip-1.42262pt(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}\hskip-1.42262pt\in\hskip-1.42262pt\mathcal{RH}_{\infty}^{2n\times 2n}$ (cf.
+proof of Lemma 4), and, given there are no RHP pole-zero cancellations between $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$, we conclude stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}\in\mathcal{RH}_{\infty}^{4n\times 4n}$.
 5 Numerical Case Studies
+To validate our theoretical stability guarantees, we conduct numerical case studies in MATLAB/Simulink using a simple three-bus system with three GFM VSCs.
+This deliberately minimal setup enables a clear interpretation and verification of the theoretical results without the added complexity of large-scale systems.
+Specifically, it allows a comparison between (i) an analytical model derived from ideal block-diagram representations of the network and device dynamics, used to obtain the stability certificates (Case Study I), and (ii) a nonlinear electromagnetic transient (EMT) simulation with more realistic models (Case Study II).
+Since the proposed stability certificates are fully decentralized, they are independent of network size and the number of units.
+Hence, while we focus on a small system for clarity, the results can be readily extended to larger systems, a direction we leave for future work.
 5.1 Case Study I: Linearized & reduced models
-We numerically validate Theorem 2 and Corollary 1 with MATLAB/Simulink using linearized small-signal models of a three-node system with three GFM VSCs.
+We numerically validate Theorem 2 and Corollary 1 using linearized small-signal models of a three-bus system with three GFM VSCs, as shown in Fig.
+10 with parameters in Table 2(b).
 We consider a network with uniform resistance-inductance ratio $\rho=0.05$, a maximum steady-state voltage magnitude $|v|_{\mathrm{max}}=1.1$ pu, and identical self susceptances $\sum_{i\neq j}^{3}b_{ij}=5$ pu of all VSC nodes $i\in\{1,2,3\}$.
-The simulations follow the ideal (i.e., linearized and reduced) block diagram dynamics in Fig.
+The simulations follow the ideal (i.e., linearized and reduced) block-diagram dynamics in Fig.
 4.
-The nodes GFM 1 and GFM 2 employ fixed controllers that always satisfy the conditions in 24 and 25, while for node GFM 3 we are exploring varying control parameters.
-We begin by modeling the network dynamics $\mathcal{N}_{0}(s)$ using Network-Simplification Level 2 in 19, without imposing additional conditions on the controller of GFM 3 (see “no cond.” in Fig.
-9).
+GFM 1 and GFM 2 employ fixed controllers that always satisfy the conditions in 25 and 26, while for GFM 3 we are exploring varying control parameters.
+[…]
+Figure 10: Schematic one-line diagram of the three-bus power system grid topology used in numerical case studies.
+Table 2: Network and GFM parameters
+(a) Electrical parameters & set points.
+Parameter Symbol Value
+Base power, voltage, frequency $S_{\mathrm{b}}$, $V_{\mathrm{b}}$, $f_{\mathrm{b}}$ 100 MW, 230 kV, 50 Hz
+RLRL line components $r_{ij}$, $l_{ij}$ 0.02, 0.4 pu
+Shunt capacitors $c_{ij}$ 0.1 pu
+Constant impedance loads $p_{\mathrm{load},i}$ 0.1 pu
+Converter power ratings $S_{1}$, $S_{2}$, $S_{3}$ 100, 100, 200 MW
+Converter $RL$ filter components $r_{\mathrm{f},i}$, $l_{\mathrm{f},i}$ 0.01, 0.1 pu
+Converter steady-state voltages $|v|_{0,1}$$|v|_{0,2}$$|v|_{0,3}$ 1, 0.9, 1.1 pu
+Converter power set points $p_{0,1}$, $p_{0,2}$, $p_{0,3}$ 0.1025, 0.1026, 0.0526 pu
+Converter power set points $q_{0,1}$, $q_{0,2}$, $q_{0,3}$ 0, -0.635, 0.3495 pu
+(b) GFM converter parameters.
+Parameter Symbol Value
+Droop gains GFM 1 (fixed) $d_{\mathrm{p,1}}$, $d_{\mathrm{q,1}}$ 0.003 pu,0.01 pu
+Time constants GFM 1 (fixed) $\tau_{\mathrm{p,1}}$, $\tau_{\mathrm{q,1}}$ 0.1 s, 0.1 s
+Droop gains GFM 2 (fixed) $d_{\mathrm{p,2}}$, $d_{\mathrm{q,2}}$ 0.003 pu,0.01 pu
+Time constants GFM 2 (fixed) $\tau_{\mathrm{p,2}}$, $\tau_{\mathrm{q,2}}$ 0.1 s, 0.1 s
+Droop gains GFM 3 (no cond.) $d_{\mathrm{p,3}}$, $d_{\mathrm{q,3}}$ 0.25 pu, 0.25 pu
+Time constants GFM 3 (no cond.) $\tau_{\mathrm{p,3}}$, $\tau_{\mathrm{q,3}}$ 0 s, 0 s
+Droop gains GFM 3 (cond. L1) $d_{\mathrm{p,3}}$, $d_{\mathrm{q,3}}$ 0.2 pu, 0.2 pu
+Time constants GFM 3 (cond. L1) $\tau_{\mathrm{p,3}}$, $\tau_{\mathrm{q,3}}$ 0.01 s, 0.01 s
+Droop gains GFM 3 (cond. DYN) $d_{\mathrm{p,3}}$, $d_{\mathrm{q,3}}$ 0.006 pu, 0.02 pu
+Time constants GFM 3 (cond. DYN) $\tau_{\mathrm{p,3}}$, $\tau_{\mathrm{q,3}}$ 0.1 s, 0.1 s
+We begin by modeling the network dynamics $\mathcal{N}_{0}(s)$ using Network-Simplification Level 2 in 20, without imposing additional conditions on the controller of GFM 3 (see “no cond.” in Fig.
+11).
 A small-signal load disturbance at node 1 reveals that the closed-loop system remains stable (Fig.
-10), confirming the validity of Corollary 1.
-Likewise, when modeling $\mathcal{N}_{0}(s)$ with Network-Simplification Level 1 in 18 and ensuring that the controller of GFM 3 satisfies the stability conditions in 26 (see “cond.
+12), confirming the validity of Corollary 1.
+Likewise, when modeling $\mathcal{N}_{0}(s)$ with Network-Simplification Level 1 in 19 and ensuring that the controller of GFM 3 satisfies the stability conditions in 27 (see “cond.
 L1” in Fig.
-9), we again observe closed-loop stability (Fig.
-11).
+11), we again observe closed-loop stability (Fig.
+13).
 This further confirms Corollary 1.
 Our main result, Theorem 2, is validated in Fig.
-12, where we model $\mathcal{N}_{0}(s)$ dynamically as in 17, and equip GFM 3 with a controller that meets the stability conditions in 24 and 25 (see “cond.
+14, where we model $\mathcal{N}_{0}(s)$ dynamically as in 18, and equip GFM 3 with a controller that meets the stability conditions in 25 and 26 (see “cond.
 DYN” in Fig.
-9).
+11).
 Stability is immediately evident.
-In contrast, using a GFM 3 controller that satisfies only 26 (“cond.
+In contrast, using a GFM 3 controller that satisfies only 27 (“cond.
 L1”) or no conditions at all (“no cond.”) leads to instability in the dynamic network model.
 This underscores the importance of accurate network modeling in control design, which is overlooked in the overly optimistic stability assessment of [8, 10, 11].
+(a) Stability conditions in 25.
+(b) Stability conditions in 26.
+Figure 11: 2D plot of the stability conditions for the three-bus system with $\rho=0.05$. The stars indicate the feasibility of the GFM 3 control parameters.
+Figure 12: Stable system response of the block diagram in Fig.
+4 with three GFM devices and the simplified network model 20, where the controller of GFM 3 does not satisfy any of the conditions 25, 26 and 27 (no cond.).
+Figure 13: Stable system response of the block diagram in Fig. 4 with three GFM devices and the simplified network model 19, where the GFM 3 controller satisfies the conditions in 27 (cond. L1).
 5.2 Case Study II: Nonlinear circuit model
-For the same three-node system as in Section 5.1, we now use Simscape Electrical in MATLAB/Simulink to perform a detailed electromagnetic transients (EMT) simulation based on a nonlinear circuit model of the three-phase network.
+For the same three-bus system as in Section 5.1, we now use Simscape Electrical in MATLAB/Simulink to perform a detailed EMT simulation based on a nonlinear circuit model of the three-phase network.
+In particular, we now consider more general and realistic line models also including shunt capacitors.
 Additionally, for each GFM VSC, we incorporate the full nonlinear converter models as shown in Fig.
-3.
-Compared to Case Study I, this increases complexity in two key aspects: first, by considering nonlinear models, and second, by accounting for the full network dynamics without the simplifying assumptions introduced in Section 3.1.
-The simulation results in Fig.
-13 demonstrate that a GFM 3 controller satisfying the decentralized stability conditions in 24 and 25 ensures stability even in the presence of nonlinear network and device models.
+3 (using average models).
+Compared to Case Study I, this increases complexity in two key aspects: first, by considering nonlinear models, and second, by accounting for the full network dynamics without the simplifying assumptions in Section 3.1, while also including shunt capacitors.
+The simulation results for a small-signal load increase in Fig.
+15 demonstrate that a GFM3 controller satisfying the decentralized stability conditions in 25 and 26 maintains stability even in the presence of nonlinear network and device models.
 Figure 14: Stable system response of the block diagram in Fig.
-4 with three GFM devices and the simplified network model 19, where the controller of GFM 3 does not satisfy any of the conditions 24, 25 and 26 (no cond.).
-Figure 11: Stable system response of the block diagram in Fig. 4 with three GFM devices and the simplified network model 18, where the GFM 3 controller satisfies the conditions in 26 (cond. L1).
-Figure 12: Stable system response of the block diagram in Fig.
-4 with three GFM devices and the dynamic network model 17, where all GFM controllers are satisfying the conditions in 24 and 25 (cond.
+4 with three GFM devices and the dynamic network model 18, where all GFM controllers are satisfying the conditions in 25 and 26 (cond.
 DYN).
-We also display the unstable dynamics of GFM 3 when the controller is satisfying no conditions (no cond.), and when it is satisfying the conditions in 26 (cond.
+We also display the unstable dynamics of GFM 3 when the controller is satisfying no conditions (no cond.), and when it is satisfying the conditions in 27 (cond.
 L1).
-Figure 13: Stable system response of the three-node converter system with nonlinear circuit and device models where all GFM controllers are satisfying the conditions in 24 and 25 (cond.
+In contrast, when the GFM 3 controller satisfies only 27 or does not meet any stability conditions, interactions with the network dynamics can lead to instability.
+We thus conclude that our stability conditions in 25 and 26, which account for network dynamics, remain effective in a nonlinear circuit scenario including shunt capacitors, provided the system operates near the nominal point where linearization errors are small.
+However, the stability conditions, designed for a simplified static network, may fail.
+Figure 15: Stable system response of the three-bus converter system with nonlinear circuit and device models (including shunt capacitors) where all GFM controllers are satisfying the conditions in 25 and 26 (cond.
 DYN).
-Additionally, we indicate the unstable dynamics of GFM 3 when the controller is satisfying no conditions (no cond.), and when it is satisfying the conditions in 26 (cond.
+Additionally, we indicate the unstable dynamics of GFM 3 when the controller is satisfying no conditions (no cond.), and the conditions in 27 (cond.
 L1).
-In contrast, when the GFM 3 controller satisfies only 26 or does not meet any stability conditions, interactions with the network dynamics can lead to instability.
-Overall, we conclude that our stability conditions in 24 and 25, which explicitly account for network dynamics, remain effective in a nonlinear circuit scenario, provided the system operates near the nominal point where linearization errors are sufficiently small.
-However, the stability conditions, designed for a simplified static network, may fail when network dynamics are taken into account.
 6 Conclusion
@@
 Our numerical case studies validated our theoretical results even for non-linear circuit and device models.
-Future work includes the extension of our dynamic loop-shifting framework to a non-diagonal structure of the matrix blocks $\Gamma_{i}(s)$ with cross-coupling between active and reactive power, thereby eventually reducing conservatism by requesting coupled device controllers.
+Future work includes the extension of our framework to a non-diagonal structure of the matrix blocks $\Gamma_{i}(s)$, thereby eventually reducing conservatism by requesting coupled device controllers.
 Beyond that, from an application point of view, we envision our stability framework to support the formulation of new grid codes for future power systems.